### v1

# Learning tensor networks with parameter dependence for Fourier-based option pricing

Abstract
A long-standing issue in mathematical finance is the speed-up of pricing options, especially multi-asset options. A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor networks to compress high-dimensional tensors. Another usage of the tensor network is to compress functions, including their parameter dependence. In this study, we propose a pricing method, where, by a tensor learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters. As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices. We show that, in the tested cases involving up to about 10 assets, the proposed method is comparable to or outperforms Monte Carlo simulation with $10^{5}$ paths in terms of computational complexity, keeping the comparable accuracy.
I Introduction
Financial firms are conducting demanding numerical calculations in their business. One of the most prominent ones is option pricing. An option is a financial contract in which one party, upon specific conditions being met, pays an amount (payoff) determined by the prices of underlying assets such as stocks and bonds to the other party. Concerning the time when the payoff occurs, the simplest and most common type of options is the European type, which this paper hereafter focuses on: at the predetermined future time (maturity) $T$, the payoff $v(\vec{S}(T))$ depending on the underlying asset prices $\vec{S}(T)$ at time $T$ occurs. For example, in a European call/put option, one party has the right to buy/sell an asset at the predetermined price (strike) $K$ and maturity $T$, and the corresponding payoff function is $v(S(T))=\max\{c(S(T)-K),0\}$, where $c=1$ and $-1$ for a call and put option, respectively. In addition to this simple one, various types of options are traded and constitute a large part of the financial industry.
Pricing options appropriately is needed for making a profit and managing the risk of loss in option trading. According to the theory in mathematical finance  [^1] As typical textbooks in this area, we refer to [1, 2]., the price of an option is given by the expectation of the discounted payoff in the contract with some stochastic model on the dynamics of underlying asset prices assumed. Except for limited cases with simple contract conditions and models, the analytical formula for the option price is not available and thus we need to resort to the numerical calculation. In the rapidly changing financial market, quick and accurate pricing is vital in option trading, but it is a challenging task, for which long-lasting research has been made. In particular, pricing multi-asset options, whose payoff depends on the prices of multiple underlying assets, is often demanding. Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
Motivated by these points, recently, applications of quantum computing to option pricing are considered actively [^2] See [3] as a comprehensive review.. For example, many studies have focused on applications of the quantum algorithm for Monte Carlo integration [4], which provides the quadratic quantum speed-up over the classical counterpart. Unfortunately, running such a quantum algorithm requires fault-tolerant quantum computers, which may take decades to be developed.
In light of this, applications of quantum-inspired algorithms to option pricing have been also studied, as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews. and recently, is utilized also in machine learning [14, 15]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [16, 17, 18, 19] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
In this study, we aim to make FT-based option pricing more efficient using TTs in a more extended way. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We then use these TTs to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to about ten assets, the computational complexity of our proposed method, which is measured by the number of elementary operations, is either advantageous or comparable to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. When the computational complexity is comparable, our numerical results indicate that our method is faster in terms of computation time, since our method involves only elementary arithmetic contrary to the Monte Carlo method involving more complicated operations. We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
The rest of this paper is organized as follows. Sec. II is devoted to introducing the tensor train and tensor learning algorithm. In Sec. III, we review the FT-based option pricing and its reformulation aided by TCI. We propose a new scheme for fast option pricing by learning tensor networks with parameter dependence in Sec. IV. We show the results of the numerical demonstration of our method applied to a kind of multi-asset option for the various values of the volatilities and the present asset prices in Sec. V. The summary and discussions are given in Sec. VI.
II Tensor train
A $d$-way tensor $F_{x_{1},\dots,x_{d}}$, where each local index $x_{l}$, $l=1,\dots,d$, has a local dimension $N$, can be decomposed into a TT format with a low-rank structure. The TT decomposition of $F_{x_{1},\ldots,x_{d}}$ can be expressed as follows.
$F_{x_{1},\ldots,x_{d}}$ $\approx\sum_{l_{1}}^{\chi_{1}}\cdots\sum_{l_{d-1}}^{\chi_{d}}F_{l_{1},x_{1}}^{(1)}F_{l_{1}l_{2},x_{2}}^{(2)}\cdots F_{l_{d-1},x_{d}}^{(d)}$
$\equiv F_{x_{1}}^{(1)}\cdot F_{x_{2}}^{(2)}\cdot(\cdots)\cdot F_{x_{d}}^{(d)}$ (1)
where $F_{x_{i}}^{(i)}$ denotes each 3-way tensor, $l_{i}$ represents the virtual bond index, and $\chi_{i}$ is the dimension of the virtual bond. One of the main advantages of TT is that it significantly reduces computational complexity and memory requirements by reducing bond dimensions $\chi_{i}$.
This is an equivalent expression to the wave function $F_{x_{1},\ldots,x_{d}}$ of a quantum system with $d$ $N$-level qudits as follows:
|$\ket{F_{x_{1},x_{2},\ldots,x_{d}}}$
$=\sum_{l_{1}}^{\chi_{1}}\cdots\sum_{l_{d-1}}^{\chi_{d}}F_{l_{1},x_{1}}^{(1)}F_{l_{1}l_{2},x_{2}}^{(2)}\cdots F_{l_{d-1},x_{d}}^{(d)}\ket{\vec{x}},$ (2)
[…]
II.1 Compression tenchniques
We introduce the two compression techniques used in this study.
II.1.1 Tensor cross interpolation
Tensor cross interpolation (TCI) is a technique to compress tensors corresponding to discretized multivariate functions with a low-rank TT representation. Here, we consider a tensor that, with grid points set in $\mathbb{R}^{d}$, has entries $F_{x_{1},\ldots,x_{d}}$ equal to $F(x_{1},\ldots,x_{d})$, the values of a function $F$ on the grid points [^4] Although we here denote the indexed of the tensor and the variables of the function by the same symbols $x_{1},\ldots,x_{d}$ for illustrative presentation, we assume that, in reality, the grid points in $\mathbb{R}^{d}$ is labeled by integers and the indexes of the tensor denotes the integers.. Leaving the detailed explanation to Refs. [16, 17, 18, 19], we describe its outline. It learns a TT using the values of the target function $F_{x_{1},x_{2}\ldots,x_{d}}$ at indexes $(x_{1},x_{2},\cdots,x_{d})$ adaptively sampled according to the specific rules. TCI actively inserts adaptively chosen interpolation points (pivots) from the sample points to learn the TT, which can be seen as a type of active learning. It gives the estimated values of the function at points across the entire domain although we use only the function values at a small number of sample points in learning. This is the very advantage of TCI and is particularly useful for compressing target tensors with a vast number of elements, contrary to singular value decomposition (SVD) requiring access to the full tensor. Note that TCI is a heuristic method, which means its effectiveness heavily depends on the internal algorithm to choose the pivots and the initial set of points selected randomly.
In this study, when we learn TT from functions with TCI, we add the pivots so that the error in the maximum norm $(\epsilon_{\mathrm{TCI}})$ is to be minimized.
$\epsilon_{\mathrm{max}}=\frac{\|F_{\mathrm{TT}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{max}}}{\|F_{\mathrm{TT}}\|_{\mathrm{max}}}$ (3)
where $F_{\mathrm{TT}}$ is a target tensor, $\tilde{F}_{\mathrm{TT}}$ is a low-rank approximation, and the maximum norm is evaluated as the maximum of the absolute values of the entries at the pivots selected already. The computational complexity of TCI is roughly proportional to the number of elements in the TT, which is $O(d\chi^{2}N)$ with $\chi_{1},\cdots,\chi_{d}$ fixed to $\chi$. In addition, considering the case that zero is included in the reference function value, the error should be normalized by $\|F_{\mathrm{TT}}\|_{\mathrm{max}}$.
II.1.2 Singular value decomposition
In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low. The original TT, denoted as $F_{\mathrm{TT}}$, can be approximated by another TT, $\tilde{F}_{\mathrm{TT}}$, with reduced (optimized) bond dimensions such that
$\frac{\|F_{\mathrm{TT}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{F}}^{2}}{\|F_{\mathrm{TT}}\|_{\mathrm{F}}^{2}}<\epsilon_{\mathrm{SVD}},$ (4)
where $\|\cdots\|_{\mathrm{F}}$ denotes the Frobenius norm. Note that the approximation error is conventionally expressed as the squared deviation. Given a tolerance $\epsilon_{\mathrm{SVD}}$, truncation via SVD results in an optimally low-rank TT approximation characterized by the smallest possible rank. For more technical details, readers are referred to Ref. [20].
III Fourier transform-based option pricing aided by tensor cross interpolation
III.1 Fourier transform-based option pricing
In this paper, we consider the underlying asset prices $\vec{S}(t)=(S_{1}(t),\cdots,S_{d}(t))$ in the Black-Scholes (BS) model described by the following stochastic differential equation
$dS_{j}(t)=rS_{j}(t)dt+\sigma_{j}S_{j}(t)dW_{j}(t).$ (5)
Here, $W_{1}(t),\cdots,W_{d}(t)$ are the Brownian motions with constant correlation matrix $\rho_{ij}$, namely
$dW_{i}(t)dW_{j}(t)=\rho_{ij}dt,$ (10)
and $r\in\mathbb{R}$ and $\sigma_{1},\cdots,\sigma_{d}>0$ are constant parameters called the risk-free interest rate and the volatilities, respectively. The present time is set to $t=0$ and the present asset prices are denoted by $\vec{S}_{0}=(S_{1,0},\cdots,S_{d,0})$.
We consider European-type options, in which the payoff $v(\vec{S}(T))$ depending on the asset prices $\vec{S}(T)$ at the maturity $T$ occurs at $T$. According to the theory of option pricing, the price $V$ of such an option is given by the expectation of the discounted payoff:
$V(\vec{p})$ $=\mathbb{E}\left[e^{-rT}v(\vec{S}(T))\middle|\vec{S}_{0}\right]$
[…]
where we define $\exp(\vec{x}):=(e^{x_{1}},\cdots,e^{x_{d}})$. $p(\vec{x}|\vec{x}_{0})$ is the probability density function of $\vec{x}:=(\log S_{1}(T),\cdots,\log S_{d}(T))$, the log asset prices at $T$, conditioned on the present value $\vec{x}_{0}=(\log S_{1,0},\cdots,\log S_{d,0})$. In the BS model defined by (5), $p(\vec{x}|\vec{x}_{0})$ is given by the $d$-variate normal distribution:
$p(\vec{x}|\vec{x}_{0})=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\exp\left(-\frac{1}{2}\left(\vec{x}-\vec{\mu}\right)^{T}\Sigma^{-1}\left(\vec{x}-\vec{\mu}\right)\right),$ (7)
where $\Sigma:=(\sigma_{i}\sigma_{j}\rho_{ij}T)_{ij}$ is the covariance matrix of $\vec{x}$ and $\vec{\mu}:=\vec{x}_{0}+\left(rT-\frac{1}{2}\sigma_{1}^{2}T,\cdots,rT-\frac{1}{2}\sigma_{d}^{2}T\right)$. Note that, in Eq. (6), we denote the option price by $V(\vec{p})$, indicating its dependence on the parameter $\vec{p}$ such as the volatilities $\vec{\sigma}=(\sigma_{1},\cdots,\sigma_{d})$ and the present asset prices $\vec{S}_{0}$.
In FT-based option pricing, we rewrite the formula (6) as the integral in the Fourier space:
[…]
Here, $\vec{z}=(z_{1},\cdots,z_{d})$ is the wavenumber vector corresponding to $\vec{x}$.
[…]
is the characteristic function, and in the BS model, it is given by
$\phi(\vec{z})=\exp\left(i\sum_{j=1}^{d}z_{j}\mu_{j}-\frac{T}{2}\sum_{j=1}^{d}\sum_{k=1}^{d}\sigma_{j}\sigma_{k}z_{j}z_{k}\rho_{jk}\right).$ (10)
[…]
$v_{\text{min}}(\vec{S}_{T})=\max\{\min\{S_{1}(T),\ldots,S_{d}(T)\}-K,0\}$ (11)
and its Fourier transform is [21]
$\hat{v}_{\min}(\vec{z})=-\frac{K^{1+i\sum_{j=1}^{d}z_{j}}}{(-1)^{d}\left(1+i\sum_{j=1}^{d}z_{j}\right)\prod_{j=1}^{d}(iz_{j})}.$ (12)
[…]
In the numerical calculation of Eq. (8), we approximate it by discretization:
$V(\vec{p})$
$=\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\phi(-\vec{z}_{\mathrm{gr},\vec{j}}-i\vec{\alpha})\hat{v}(\vec{z}_{\mathrm{gr},\vec{j}}+i\vec{\alpha})\Delta_{\rm vol}.$ (13)
Here, the even natural number $N$ is the number of the grid points in one dimension. $\vec{z}_{\mathrm{gr},\vec{j}}$ is the grid point specified by the integer vector $\vec{j}$ as
$\vec{z}_{\mathrm{gr},\vec{j}}:=(\eta_{1}(\vec{p})j_{1},\ldots,\eta_{d}(\vec{p})j_{d}),$ (14)
where $\eta_{j}$ is the integration step size in the $j$-th direction. $\eta_{j}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{j}(\vec{p})$ in (14). In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec. V for the detail. $\Delta_{\rm vol}:=\Pi_{j=1}^{d}\eta_{j}$ is the volume element.
III.2 Fourier transform-based option pricing with tensor trains
Note that to compute the sum in Eq. (13), we need to evaluate $\phi$ and $\hat{v}$ exponentially many times with respect to the asset number $d$. This is not feasible for large $d$. Then, to reduce the computational complexity, following Ref. [6], we consider approximating $\phi$ and $\hat{v}$ by TTs. For the tensor $\phi_{j_{1},\ldots,j_{d}}$ (resp. $\hat{v}_{j_{1},\ldots,j_{d}}$), whose entry with index $\vec{j}$ is $\phi(-\vec{z}_{\mathrm{gr},\vec{j}}-i\vec{\alpha})$ (resp. $\hat{v}(\vec{z}_{\mathrm{gr},\vec{j}}+i\vec{\alpha})$), we construct a TT approximation $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ (resp. $\tilde{v}_{j_{1},\ldots,j_{d}}$) by TCI. Then, we approximately calculate Eq. (13) by
$V(\vec{\sigma},\vec{S}_{0})\simeq\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\tilde{\phi}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}}.$ (15)
Thanks to TCI, we can obtain the approximate TTs avoiding the evaluations of $\phi$ and $\hat{v}$ at all the grid points. Besides, given the TTs, we can compute the sum in (15) as the contraction of two TTs without exponentially many iterations: with the bond dimensions at most $\chi$, the number of multiplications and additions is of order $O(d\chi^{3})$.
Hereafter, we simply call this approach for FT-based option pricing aided by TTs TT-based option pricing.
III.3 Monte Carlo-based option pricing
Here, we also make a brief description of the Monte Carlo (MC)-based option pricing. It is a widely used approach in practice, and we take it as a comparison target in our numerical demonstration of TT-based option pricing.
In the MC-based approach, we estimate the expectation in Eq. (6) by the average of the payoffs in the sample paths:
$V(\vec{p})\approx e^{-rT}\times\frac{1}{N_{\rm path}}\sum_{i=1}^{N_{\rm path}}v\left(\exp(\vec{x}_{i})\right),$ (16)
where $\vec{x}_{1},\ldots,\vec{x}_{N_{\rm path}}$ are i.i.d. samples from $p(\vec{x}|\vec{x}_{0})$. On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., [22]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions. Besides, calculating the payoff $v$ with the normal variable $\vec{x}_{i}$ involves exponentiation. In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
IV Learning tensor trains with parameter dependence
IV.1 Outline
Option prices depend on input parameters such as the volatilities $\vec{\sigma}$ and the present asset prices $\vec{S}_{0}$. In the rapidly changing financial market, these input parameters vary from time to time, which causes the change of the option price. Therefore, if we have a function that efficiently outputs an accurate approximation of the option price for various values of the input parameters, it provides a large benefit to practical business.
Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set. Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions. The outline is illustrated in Fig. 1. Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$. Then, as illustrated in Fig. 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$. Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$. Having these tensor trains, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig. 1 (b). By fixing the index $\vec{k}$ of $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ to the value corresponding to the specified $\vec{p}$ and contracting the adjacent core tensors pairwise, we get the $d$-dimensional TTs $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$. Contracting these TTs as Eq. (15), we get the option price for the specified $\vec{p}$.
Here, we note that the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ for $\hat{v}$ and its TT approximation $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ corresponding to $\vec{p}$, which is now $\vec{\sigma}$ or $\vec{S}_{0}$, the parameter not affecting $\hat{v}$. This is due to $\eta_{j}$ depending on $\vec{p}$: because of the change of the grid points in the $\vec{z}$ space with respect to $\vec{p}$, the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ depends on the indexes $k_{i}$.
We also mention the ordering of the local indices of the TT. Two core tensors in the TT that correspond to $z_{j}$ and $p_{j}$ of the same asset are arranged next to each other, i.e., the order is ($z_{1}p_{1}z_{2}p_{2}\cdots z_{d}p_{d}$). In the demonstration in Sec. V, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing. On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains. However, the optimal arrangement of the local indices may vary, for example, depending on the correlation matrix: intuitively, the core tensors corresponding to highly correlated assets should be placed nearby. Although we do not discuss it in detail, this is an important topic for future research.
In the two test cases for our proposed method in Sec. V, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices. Note that it is possible to include dependencies on other parameters in the TTs. For example, taking into account the dependence on the parameters concerning the option contract, such as the maturity $T$ and the strike $K$, enables us to price different option contracts with a single set of TTs. Although this is a promising approach, there might be some issues. For example, it is non-trivial whether TTs incorporating many parameter dependencies have a low-rank structure. Thus, we will leave such a study for future work.
Figure 1: Fast option pricing based on TTs proposed in this study. In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD. In (b), we use these TTs to perform fast option pricing for a specified parameter $\vec{p}$.
IV.2 Computational complexity
The computational complexity of TT-based option pricing, involving the contraction of tensors associated with $p_{i}$ and $z_{i}$ and the computation of the inner product between two TTs, is $O(d\chi^{3}N)$ for both computations. In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of learning the TTs by TCI and SVD and consider that after we get the TTs. This is reasonable if we can use plenty of time to learn the TTs before we need fast option pricing. As discussed in Sec. VI, we can reasonably find such a situation in practice.
V Numerical demonstration
V.1 Details
Now, as a demonstration, we apply the proposed method to pricing a $d$-asset European min-call option in the BS model. In the following, we describe the parameter values used in this study, the software used, and how the errors were evaluated.
Ranges of ${\sigma_{i}}$ and ${S^{i}_{0}}$
With respect to $\vec{p}$, on which the TTs learn the dependence of the functions $\phi$ and $\hat{v}$, we take the two test cases: $\vec{p}$ is the volatilities $\vec{\sigma}$ or the present asset prices $\vec{S}_{0}$. In the proposed method, we need to set the range in the space of $\vec{p}$ and the grid points in it. For each volatility $\sigma_{i}$, we set the range to $\sigma_{i}\in[0.15,0.25]$, where the center $\sigma_{i}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [23] and the range width ±$\pm 0.05$ covers the changes in this index on most days. For each $S_{i,0}$, we set the range to $S_{i,0}\in[90,120]$, which corresponds to the 20% variation of the asset price centered at 100. The lower bound is set to not 80 but 90 because the price of the option we take as an example is negligibly small for $S_{0}<90$. For both $\sigma_{i}$ and $S_{i,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
Other parameters
The other parameters for option pricing are fixed to the values summarized in the table below.
$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{ii}$ $\rho_{ij(i\neq j)}$ $N$
$1$ $0.01$ $100$ $100$ $5/d$ $1$ $1/3$ $50$
Table 1: The parameters except $\vec{\sigma}$ and $\vec{S}_{0}$.
Integral step size
The spread of the characteristic function changes depending on the values of the parameters $\vec{p}$ and the number of assets. Accordingly, it is necessary to adjust the grid interval $\eta_{j}$ for integration in the wavenumber space.
For example, consider the case where $\vec{p}=\vec{\sigma}$. From Eq. (10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{j}z_{j}\lesssim 1$ for each $j$. It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval. Currently, as $z_{j}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{j}$, $\eta_{j}$ and $\sigma_{j}$ must satisfy $\sigma_{j}\eta_{j}\ll 1$ and $\sigma_{j}\times N\eta_{j}\gg 1$. Therefore, in general, we need the adjustment that $\eta_{j}$ increases when $\sigma$ decreases and vice versa. Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{j}$ set to some constant, and thus we did so. Only in the cases with $d=10$, $\eta_{j}$ is adjusted as
$\eta_{j}=\frac{a}{\sigma_{j}}+b$ (17)
To set the constants $a$ and $b$, we ran TT-based option pricing for several values of $\sigma_{j}$ adjusting $\eta_{j}$, found the pairs $(\sigma_{j},\eta_{j})$ that yields the accurate price, and adjust $a$ and $b$ so that Eq. (17) approximately reproduces such $\eta_{j}$ from $\sigma_{j}$.
Also when $\vec{p}=\vec{S}_{0}$, we set $\eta_{j}$ to some constant in almost all the test cases. Only for $d=11$, we try the adjustment that
$\eta_{j}=\frac{\alpha}{S_{j,0}}+\beta$ (18)
with $\alpha$ and $\beta$ chosen according to the option pricing result for several $\eta_{j}$ and $\sigma_{j}$. In general, larger $S_{j,0}$, which means larger $x_{j,0}$, makes $\phi$ as a function of $z_{j}$ oscillate more rapidly, and thus we need finer grid points to resolve such an oscillation.
Note that when $\eta_{j}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$. In such a case, the contraction between the core tensor for $z_{j}$ and that for $p_{j}$ is not needed since the latter does not exist, and thus the computational complexity decreases.
Error evaluation
We do not have the exact price of the multivariate min-call option since there is no known analytic formula for it. Instead, we regard the option price computed by the MC-based method with very many paths, concretely $5\times 10^{7}$ paths, as the true value. The error of the option price computed by the proposed method is evaluated by its deviation from the true value. We take the MC-based method with $10^{5}$ paths, which is a typical number in option pricing in practical business, as the comparison target against our method. The average relative error and its standard deviation of the MC-based method with $10^{5}$ paths are calculated from its 20 runs. We assess the accuracy of our method by seeing if it falls within the statistical error range of the MC-based method.
An issue is that the number of possible combinations of the parameter $\vec{p}$ is $100^{d}$, and thus we cannot test all of them. Thus, we randomly select 100 combinations and perform option pricing for each of them. We compare the maximum among the relative errors of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
Software and hardware used in this study
The TCI code has not yet been made publicly available. Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft provided us with access to their TCI code. ITensors.jl [24] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [25]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
V.2 Results
We show the results for the computational complexity, time, and accuracy of TT-based and MC-based option pricing when two parameters $\vec{\sigma}$ and $\vec{S_{0}}$ are varied.
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational times of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$, and subsequently reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{v}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
V.2.1 The case of varying $\vec{\sigma}$
Figure 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. Up to $d=9$, TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI and subsequently recompressing, the bond dimensions could generally be maintained at 10 or below. The details of this compression by SVD are described in Appendix A. Furthermore, The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
[…]
When the computational complexity of the two approaches is comparable, we argue that TT-based option pricing has an advantage. This is because, in the MC-based method, the unit operations such as random number generation are more complex than those in the TT-based method, which are just multiplications and additions. Indeed, we see that in our demonstration, the TT-based method is in fact advantageous in terms of computational time with the software we use. However, since the computational complexity of the two methods is not significantly different, we should note that the quality of the implementation could affect their computational time.
It has been observed that accuracy tends to decrease as $d$ increases. Consequently, increasing $d$ beyond 10 results in decreased accuracy and makes it difficult to maintain accuracy within that in the MC-based method. At the same time, the bond dimension increases, which makes it difficult to maintain the computational complexity smaller than the MC-based method.
Furthermore, reducing the number of subdivisions below 100 resulted in only minor changes in accuracy and computational complexity.
(a) $\vec{\sigma}$
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec]
$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $7.02\times 10^{4}$ $5.0\times 10^{5}$ $3.79\times 10^{-4}$ $2.88\times 10^{-2}$
$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $1.24\times 10^{5}$ $6.0\times 10^{5}$ $5.88\times 10^{-4}$ $3.38\times 10^{-2}$
$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0202 0.0236 $2.08\times 10^{5}$ $7.0\times 10^{5}$ $7.02\times 10^{-4}$ $3.71\times 10^{-2}$
$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0180 0.0253 $2.59\times 10^{5}$ $8.0\times 10^{5}$ $1.46\times 10^{-3}$ $4.09\times 10^{-2}$
$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $3.80\times 10^{5}$ $9.0\times 10^{5}$ $2.51\times 10^{-3}$ $4.51\times 10^{-2}$
$10$ 0.58 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0487 0.0337 $3.72\times 10^{5}$ $1.0\times 10^{6}$ $2.62\times 10^{-3}$ $4.88\times 10^{-2}$
$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0308 0.0357 $1.86\times 10^{6}$ $1.0\times 10^{6}$ $6.08\times 10^{-3}$ $5.11\times 10^{-2}$
(b) $\vec{S}_{0}$
$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec]
$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $5.63\times 10^{4}$ $5.0\times 10^{5}$ $6.58\times 10^{-4}$ $2.90\times 10^{-2}$
$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $1.16\times 10^{5}$ $6.0\times 10^{5}$ $7.92\times 10^{-4}$ $3.28\times 10^{-2}$
$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $1.78\times 10^{5}$ $7.0\times 10^{5}$ $1.15\times 10^{-3}$ $3.66\times 10^{-2}$
$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0258 0.0340 $3.06\times 10^{5}$ $8.0\times 10^{5}$ $1.67\times 10^{-3}$ $3.93\times 10^{-2}$
$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $4.27\times 10^{5}$ $9.0\times 10^{5}$ $2.98\times 10^{-3}$ $4.54\times 10^{-2}$
$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $6.36\times 10^{5}$ $1.0\times 10^{6}$ $4.47\times 10^{-3}$ $4.93\times 10^{-2}$
$11$ 0.48 $6.0\times 10^{-6}$ $6.0\times 10^{-6}$ 0.0389 0.0202 $6.38\times 10^{5}$ $1.1\times 10^{6}$ $4.22\times 10^{-3}$ $4.88\times 10^{-2}$
$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0377 0.0401 $4.83\times 10^{6}$ $1.1\times 10^{6}$ $1.73\times 10^{-2}$ $5.28\times 10^{-2}$
Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence. Here, we set the ranges $\sigma_{i}\in[0.15,0.25]$ and $S_{0}^{i}\in[90,120]$, and place 100 equally spaced grid points within these ranges. The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$. For both (a) $\vec{\sigma},d=10$ and (b) $\vec{S}_{0},d=11$, there are two scenarios for $\eta$: constant and auto. The term auto in (a) and (b) indicate that $\eta_{i}$ was automatically determined for each asset based on Eqs. (17) and (18), respectively. In the constant scenario, the maximum relative error does not stay within the error range of MC with PATH$=10^{5}$, whereas the auto scenario achieves success. $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, versus the number of assets $d$. We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs. In TT-based option pricing, the sharp increase in computational complexity around $d=10$ arises. This stems from the necessity to set $\eta$ appropriately for each asset, which in turn makes the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ for $\vec{p}$.
Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$. The odd bond $l=2i-1$ connects the core tensors for $j_{i}$ and $k_{i}$, the indexes for $z_{i}$ and $p_{i}$, respectively, and the even bond $l=2i$ connects those for $k_{i}$ and $j_{i+1}$. In the cases corresponding to the left panels and the bottom right panel, $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{i}$ and $z_{i+1}$). It is noteworthy that the graph of bond dimensions exhibits a characteristic jagged shape. The bond dimensions between sites for $j_{i}$ and $k_{i}$, which are related to the same asset, are large, and the ones between sites for $k_{i}$ and $j_{i+1}$, which are related to different assets, are small.
V.2.2 The case of varying $\vec{S_{0}}$
Figure 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence. Up to $d=10$, TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time. At $d$=5, after compression using SVD, the bond dimensions were reduced to around 10 (refer to Fig. 3(b)). However, at $d=11$, due to the same reasons as in the case of $\sigma$, the computational complexity of TT-based calculations becomes comparable to that of the Monte Carlo method. Nonetheless, due to the different units of computational complexity, we saw superiority over the Monte Carlo method in terms of computation time. Compared to varying $\sigma$, its accuracy is worse at $d=10$.
V.2.3 Randomness in learning the TTs
Here, we mention the randomness in learning the TTs and the error induced by it. Since TCI starts with the randomly selected grid points, the accuracy of the TT-based method depends on these initial points. To assess such a fluctuation of the accuracy, for $d=10$, we evaluated the mean and standard deviation of the relative accuracy of the TT-based method in 20 runs, with initial points randomly selected in each. As a result, the mean was 0.0324 and the standard deviation was 0.0039, which indicates that the accuracy fluctuation of TT-based option pricing is very small. We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well, that is, the error defined by Eq. (3) might not go below the threshold. Here, with the tolerance for TCI/SVD fixed, if the maximum relative error exceeds 4%, we redo TCI and SVD. The aforementioned mean and standard deviation are computed from successful cases.
VI Summary and discussion
We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs. (10) and (12), enabling fast option pricing in response to varying parameters. In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method. Up to $d=9$, we demonstrated superiority in both computational complexity and time. For $d=10$ or $11$, the proposed method is comparable to the Monte Carlo method from the perspective of computational complexity. This is partly due to the need to adjust the integration mesh width $\eta$ for each asset, resulting in an increase in bond dimensions and, consequently, the computational complexity of contraction. However, we argue that if the computational complexity is comparable, TT-based option pricing holds an advantage in terms of computation time. This is because the TT-based method involves only multiplications and additions as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, iterated $dN_{\rm path}$ times. When the computational complexity of both methods was comparable, we confirmed that our proposed method was advantageous in terms of computation time. At this point, we also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths. However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\simeq 10$.
Now, let us consider how the proposed method provides benefits in practical business in financial firms. An expected way to utilize this method is as follows. At night when the financial market is closed, we learn the TTs, and then, during the day when the market is open, we use these TTs to quickly price the option for the fluctuating parameters. If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial. In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process.
Finally, we discuss future research directions. If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing. It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g. by increasing $N$ from the current value of 50 to, say, 100. It is a promising approach to incorporate parameter dependence using the Matrix Product Operators (MPO). It could eliminate the contraction between tensors related to parameters $p_{i}$ and $z_{i}$, potentially speeding up option pricing further. At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself. For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TT format is desired, although this study tried to incorporate either of them as the first step. Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.
Acknowledgements.
We are grateful to Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the code for tensor cross interpolation. R. S. is supported by the JSPS KAKENHI Grant No. 23KJ0295. R. S. thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor networks to option pricing. K. M. is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant no. JPMXS0120319794, JSPS KAKENHI Grant no. JP22K11924, and JST COI-NEXT Program Grant No. JPMJPF2014.
Appendix A How to set the tolerance of SVD
Figure 4: The maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with various $\epsilon_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$. The left vertical axis represents $r_{\mathrm{TT}}$, while the right vertical axis shows $c_{\mathrm{TT}}$.
Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI. The cutoff $\epsilon_{\text{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$. From Fig. 4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ while minimizing the computational complexity.
From the fact that we can keep the maximum relative error $r_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information. By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed. In addition, it is surprising that in the analysis with $\epsilon_{\text{SVD}}=10^{-6}$, the maximum relative error decreased compared to before compression by SVD. We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance. We expect that this phenomenon does not occur generally, and in fact, it did not occur for other asset numbers or parameters.

### v2

@@
-Learning tensor networks with parameter dependence for Fourier-based option pricing
+Learning parameter dependence for Fourier-based option pricing with tensor networks
 Abstract
-A long-standing issue in mathematical finance is the speed-up of pricing options, especially multi-asset options.
+A long-standing issue in mathematical finance is the speed-up of option pricing, especially for multi-asset options.
 A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor networks to compress high-dimensional tensors.
 Another usage of the tensor network is to compress functions, including their parameter dependence.
-In this study, we propose a pricing method, where, by a tensor learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters.
+Here, we propose a pricing method, where, by a tensor learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters.
 As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices.
-We show that, in the tested cases involving up to about 10 assets, the proposed method is comparable to or outperforms Monte Carlo simulation with $10^{5}$ paths in terms of computational complexity, keeping the comparable accuracy.
+We show that, in the tested cases involving up to 11 assets, the proposed method outperforms MC-based option pricing with $10^{5}$ paths in terms of computational complexity while keeping comparable accuracy.
 I Introduction
@@
 On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews.
-and recently, is utilized also in machine learning [14, 15].
+and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17], and differential equations [7, 18].
 Ref.
@@
 The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [19, 20, 17, 21] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
-In this study, we aim to make FT-based option pricing more efficient using TTs in a more extended way.
+Here, we would like to point out an issue in this TT-based method.
+That is, we need to rerun TCI to obtain TTs each time the parameters are changed.
+According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
+To address these issues, we focus on another usage of tensor networks that can embed parameter dependence [22, 23, 24] to make FT-based option pricing more efficient.
 Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function.
-We then use these TTs to perform fast option pricing in response to various parameter changes.
+We then use these tensor networks to perform fast option pricing in response to various parameter changes.
 To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices.
-In the test cases, it is seen that for up to about ten assets, the computational complexity of our proposed method, which is measured by the number of elementary operations, is either advantageous or comparable to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice.
-When the computational complexity is comparable, our numerical results indicate that our method is faster in terms of computation time, since our method involves only elementary arithmetic contrary to the Monte Carlo method involving more complicated operations.
+In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice.
+When the computational complexity is smaller, our numerical results indicate that our method has a further advantage in terms of computation time, since our method involves only elementary arithmetic contrary to the Monte Carlo method involving more complicated operations.
 We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths.
@@
 $dW_{i}(t)dW_{j}(t)=\rho_{ij}dt,$ (10)
-and $r\in\mathbb{R}$ and $\sigma_{1},\cdots,\sigma_{d}>0$ are constant parameters called the risk-free interest rate and the volatilities, respectively.
+where $r\in\mathbb{R}$ and $\sigma_{1},\cdots,\sigma_{d}>0$ are constant parameters called the risk-free interest rate and the volatilities, respectively.
 The present time is set to $t=0$ and the present asset prices are denoted by $\vec{S}_{0}=(S_{1,0},\cdots,S_{d,0})$.
@@
 where we define $\exp(\vec{x}):=(e^{x_{1}},\cdots,e^{x_{d}})$.
-$p(\vec{x}|\vec{x}_{0})$ is the probability density function of $\vec{x}:=(\log S_{1}(T),\cdots,\log S_{d}(T))$, the log asset prices at $T$, conditioned on the present value $\vec{x}_{0}=(\log S_{1,0},\cdots,\log S_{d,0})$.
-In the BS model defined by (5), $p(\vec{x}|\vec{x}_{0})$ is given by the $d$-variate normal distribution:
-$p(\vec{x}|\vec{x}_{0})=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\exp\left(-\frac{1}{2}\left(\vec{x}-\vec{\mu}\right)^{T}\Sigma^{-1}\left(\vec{x}-\vec{\mu}\right)\right),$ (7)
+$q(\vec{x}|\vec{x}_{0})$ is the probability density function of $\vec{x}:=(\log S_{1}(T),\cdots,\log S_{d}(T))$, the log asset prices at $T$, conditioned on the present value $\vec{x}_{0}=(\log S_{1,0},\cdots,\log S_{d,0})$.
+In the BS model defined by (5), $q(\vec{x}|\vec{x}_{0})$ is given by the $d$-variate normal distribution:
+$q(\vec{x}|\vec{x}_{0})=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\exp\left(-\frac{1}{2}\left(\vec{x}-\vec{\mu}\right)^{T}\Sigma^{-1}\left(\vec{x}-\vec{\mu}\right)\right),$ (7)
 where $\Sigma:=(\sigma_{i}\sigma_{j}\rho_{ij}T)_{ij}$ is the covariance matrix of $\vec{x}$ and $\vec{\mu}:=\vec{x}_{0}+\left(rT-\frac{1}{2}\sigma_{1}^{2}T,\cdots,rT-\frac{1}{2}\sigma_{d}^{2}T\right)$.
@@
 where $\vec{x}_{1},\ldots,\vec{x}_{N_{\rm path}}$ are i.i.d.
-samples from $p(\vec{x}|\vec{x}_{0})$.
+samples from $q(\vec{x}|\vec{x}_{0})$.
 On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., [27]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions.
@@
 In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
-IV Learning tensor trains with parameter dependence
+IV Learning parameter dependence with tensor trains
 IV.1 Outline
@@
 1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$.
+We reduce the bond dimensions of these TTs using SVD.
+Subsequently, we contract adjacent core tensors pairwise and obtain matrix product operators (MPOs).
 Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$.
-Having these tensor trains, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig.
+Having these MPOs, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig.
 1 (b).
-By fixing the index $\vec{k}$ of $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ to the value corresponding to the specified $\vec{p}$ and contracting the adjacent core tensors pairwise, we get the $d$-dimensional TTs $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$.
+By fixing the index $\vec{k}$ of $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ to the value corresponding to the specified $\vec{p}$, we get the $d$-dimensional TTs $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$.
 Contracting these TTs as Eq.
@@
 In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD.
-In (b), we use these TTs to perform fast option pricing for a specified parameter $\vec{p}$.
+Then, we contract the tensors associated with $p_{i}$ and $z_{i}$, resulting in MPOs.
+In (b), we use these MPOs to perform fast option pricing for a specified parameter $\vec{p}$.
 IV.2 Computational complexity
-The computational complexity of TT-based option pricing, involving the contraction of tensors associated with $p_{i}$ and $z_{i}$ and the computation of the inner product between two TTs, is $O(d\chi^{3}N)$ for both computations.
+The computational complexity of TT-based option pricing, involving computation of the inner product between two TTs, is $O(d\chi^{3}N)$.
 In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations.
@@
 V.
-Here, we ignore the computational complexity of learning the TTs by TCI and SVD and consider that after we get the TTs.
-This is reasonable if we can use plenty of time to learn the TTs before we need fast option pricing.
+Here, we ignore the computational complexity of learning the TTs by TCI, SVD, and contraction of tensors associated with $p_{i}$ and $z_{i}$ and consider that after we get the MPOs only.
+This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing.
 As discussed in Sec.
@@
 Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{j}$ set to some constant, and thus we did so.
-Only in the cases with $d=10$, $\eta_{j}$ is adjusted as
+In the cases with $d=10,11$, $\eta_{j}$ is adjusted as
 $\eta_{j}=\frac{a}{\sigma_{j}}+b$ (17)
@@
 Note that when $\eta_{j}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.
-In such a case, the contraction between the core tensor for $z_{j}$ and that for $p_{j}$ is not needed since the latter does not exist, and thus the computational complexity decreases.
 Error evaluation
@@
 Also, the computational times of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively.
-To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$, and subsequently reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{v}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
+To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$.
+Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{v}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
+The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi^{\phi}_{\mathrm{max}}$ and $\chi^{\hat{v}}_{\mathrm{max}}$, respectively, in TABLE 2.
 V.2.1 The case of varying $\vec{\sigma}$
-Figure 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence.
-Up to $d=9$, TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method.
-The bond dimensions of the TT results for $d=5$ are depicted in Fig.
+TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence.
+TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method.
+The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in Fig.
 3(a).
-By applying SVD to the TTs trained via TCI and subsequently recompressing, the bond dimensions could generally be maintained at 10 or below.
+By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ could generally be maintained at less than 10 (refer to TABLE 2).
 The details of this compression by SVD are described in Appendix A.
 Furthermore, The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
-[…]
-When the computational complexity of the two approaches is comparable, we argue that TT-based option pricing has an advantage.
+We argue that, when the computational complexity of the two approaches is at least comparable, the TT-based option pricing has an advantage in terms of computational time.
 This is because, in the MC-based method, the unit operations such as random number generation are more complex than those in the TT-based method, which are just multiplications and additions.
-Indeed, we see that in our demonstration, the TT-based method is in fact advantageous in terms of computational time with the software we use.
-However, since the computational complexity of the two methods is not significantly different, we should note that the quality of the implementation could affect their computational time.
+Indeed, in our demonstration, we have shown that the TT-based method is advantageous in terms of computational complexity and exhibits even better performance in terms of computational time.
+However, it is important to note that the actual wall clock time may vary depending on the quality of the implementation.
 It has been observed that accuracy tends to decrease as $d$ increases.
-Consequently, increasing $d$ beyond 10 results in decreased accuracy and makes it difficult to maintain accuracy within that in the MC-based method.
-At the same time, the bond dimension increases, which makes it difficult to maintain the computational complexity smaller than the MC-based method.
+Consequently, increasing $d$ further might result in decreased accuracy, making it difficult to maintain accuracy within the MC-based method.
 Furthermore, reducing the number of subdivisions below 100 resulted in only minor changes in accuracy and computational complexity.
 (a) $\vec{\sigma}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec]
-$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $7.02\times 10^{4}$ $5.0\times 10^{5}$ $3.79\times 10^{-4}$ $2.88\times 10^{-2}$
-$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $1.24\times 10^{5}$ $6.0\times 10^{5}$ $5.88\times 10^{-4}$ $3.38\times 10^{-2}$
-$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0202 0.0236 $2.08\times 10^{5}$ $7.0\times 10^{5}$ $7.02\times 10^{-4}$ $3.71\times 10^{-2}$
-$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0180 0.0253 $2.59\times 10^{5}$ $8.0\times 10^{5}$ $1.46\times 10^{-3}$ $4.09\times 10^{-2}$
-$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $3.80\times 10^{5}$ $9.0\times 10^{5}$ $2.51\times 10^{-3}$ $4.51\times 10^{-2}$
-$10$ 0.58 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0487 0.0337 $3.72\times 10^{5}$ $1.0\times 10^{6}$ $2.62\times 10^{-3}$ $4.88\times 10^{-2}$
-$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0308 0.0357 $1.86\times 10^{6}$ $1.0\times 10^{6}$ $6.08\times 10^{-3}$ $5.11\times 10^{-2}$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
+$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $4.76\times 10^{4}$ $5.0\times 10^{5}$ $3.27\times 10^{-4}$ $2.84\times 10^{-2}$ $4$ $7$
+$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $7.94\times 10^{4}$ $6.0\times 10^{5}$ $6.11\times 10^{-4}$ $3.38\times 10^{-2}$ $5$ $7$
+$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0202 0.0236 $1.21\times 10^{5}$ $7.0\times 10^{5}$ $7.66\times 10^{-4}$ $3.71\times 10^{-2}$ $6$ $7$
+$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0180 0.0253 $1.48\times 10^{5}$ $8.0\times 10^{5}$ $9.07\times 10^{-4}$ $4.09\times 10^{-2}$ $6$ $7$
+$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $2.01\times 10^{5}$ $9.0\times 10^{5}$ $1.24\times 10^{-3}$ $4.51\times 10^{-2}$ $7$ $7$
+$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0308 0.0357 $2.48\times 10^{5}$ $1.0\times 10^{6}$ $2.19\times 10^{-3}$ $5.11\times 10^{-2}$ $7$ $7$
+$11$ auto $3.5\times 10^{-6}$ $3.5\times 10^{-6}$ 0.0341 0.0347 $3.02\times 10^{5}$ $1.1\times 10^{6}$ $2.57\times 10^{-3}$ $5.04\times 10^{-2}$ $8$ $7$
 (b) $\vec{S}_{0}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec]
-$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $5.63\times 10^{4}$ $5.0\times 10^{5}$ $6.58\times 10^{-4}$ $2.90\times 10^{-2}$
-$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $1.16\times 10^{5}$ $6.0\times 10^{5}$ $7.92\times 10^{-4}$ $3.28\times 10^{-2}$
-$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $1.78\times 10^{5}$ $7.0\times 10^{5}$ $1.15\times 10^{-3}$ $3.66\times 10^{-2}$
-$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0258 0.0340 $3.06\times 10^{5}$ $8.0\times 10^{5}$ $1.67\times 10^{-3}$ $3.93\times 10^{-2}$
-$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $4.27\times 10^{5}$ $9.0\times 10^{5}$ $2.98\times 10^{-3}$ $4.54\times 10^{-2}$
-$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $6.36\times 10^{5}$ $1.0\times 10^{6}$ $4.47\times 10^{-3}$ $4.93\times 10^{-2}$
-$11$ 0.48 $6.0\times 10^{-6}$ $6.0\times 10^{-6}$ 0.0389 0.0202 $6.38\times 10^{5}$ $1.1\times 10^{6}$ $4.22\times 10^{-3}$ $4.88\times 10^{-2}$
-$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0377 0.0401 $4.83\times 10^{6}$ $1.1\times 10^{6}$ $1.73\times 10^{-2}$ $5.28\times 10^{-2}$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
+$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $3.53\times 10^{4}$ $5.0\times 10^{5}$ $2.47\times 10^{-4}$ $2.90\times 10^{-2}$ $4$ $6$
+$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $6.38\times 10^{4}$ $6.0\times 10^{5}$ $5.61\times 10^{-4}$ $3.28\times 10^{-2}$ $5$ $6$
+$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $1.00\times 10^{5}$ $7.0\times 10^{5}$ $6.80\times 10^{-4}$ $3.66\times 10^{-2}$ $6$ $7$
+$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0258 0.0340 $1.54\times 10^{5}$ $8.0\times 10^{5}$ $9.71\times 10^{-4}$ $3.93\times 10^{-2}$ $7$ $7$
+$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $1.99\times 10^{5}$ $9.0\times 10^{5}$ $1.37\times 10^{-3}$ $4.54\times 10^{-2}$ $7$ $7$
+$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $3.00\times 10^{5}$ $1.0\times 10^{6}$ $1.88\times 10^{-3}$ $4.93\times 10^{-2}$ $7$ $8$
+$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0377 0.0401 $7.33\times 10^{5}$ $1.1\times 10^{6}$ $6.25\times 10^{-3}$ $5.28\times 10^{-2}$ $9$ $14$
 Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
@@
 The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$.
-For both (a) $\vec{\sigma},d=10$ and (b) $\vec{S}_{0},d=11$, there are two scenarios for $\eta$: constant and auto.
 The term auto in (a) and (b) indicate that $\eta_{i}$ was automatically determined for each asset based on Eqs.
 (17) and (18), respectively.
-In the constant scenario, the maximum relative error does not stay within the error range of MC with PATH$=10^{5}$, whereas the auto scenario achieves success.
 $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
@@
 We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs.
-In TT-based option pricing, the sharp increase in computational complexity around $d=10$ arises.
-This stems from the necessity to set $\eta$ appropriately for each asset, which in turn makes the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ for $\vec{p}$.
+In TT-based option pricing with $\vec{S_{0}}$, the sharp increase in computational complexity at $d=11$ arises.
+This stems from the larger bond dimensions with a sufficiently low cutoff value for SVD in order to maintain its accuracy
 Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$.
@@
 V.2.2 The case of varying $\vec{S_{0}}$
-Figure 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence.
-Up to $d=10$, TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time.
-At $d$=5, after compression using SVD, the bond dimensions were reduced to around 10 (refer to Fig.
+TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence.
+TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time.
+For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ were reduced to around 10 (refer to Fig.
 3(b)).
-However, at $d=11$, due to the same reasons as in the case of $\sigma$, the computational complexity of TT-based calculations becomes comparable to that of the Monte Carlo method.
-Nonetheless, due to the different units of computational complexity, we saw superiority over the Monte Carlo method in terms of computation time.
+However, at $d$ = 11, the computational complexity of TT-based calculation increases significantly.
+This is because it was necessary to increase the bond dimension by setting the SVD cutoff sufficiently low to maintain accuracy (refer to TABLE 2).
+This suggests that $d=11$ may already be reaching the upper limit of $d$ where accuracy can be maintained while surpassing the computational complexity over MC-based calculations.
 Compared to varying $\sigma$, its accuracy is worse at $d=10$.
@@
 As a result, the mean was 0.0324 and the standard deviation was 0.0039, which indicates that the accuracy fluctuation of TT-based option pricing is very small.
-We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well, that is, the error defined by Eq.
+We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well.
+That is, the error defined by Eq.
 (3) might not go below the threshold.
@@
 The aforementioned mean and standard deviation are computed from successful cases.
+V.2.4 Total computational time for obtaining the MPOs
+Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ through TCI and SVD.
+We focus on TCI because it dominates over SVD.
+It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved.
+This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec.
+VI.
+We also note that, when we do not incorporate the parameter dependence into the tensor networks for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence.
+In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1.
+This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
 VI Summary and discussion
@@
 In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method.
-Up to $d=9$, we demonstrated superiority in both computational complexity and time.
-For $d=10$ or $11$, the proposed method is comparable to the Monte Carlo method from the perspective of computational complexity.
-This is partly due to the need to adjust the integration mesh width $\eta$ for each asset, resulting in an increase in bond dimensions and, consequently, the computational complexity of contraction.
-However, we argue that if the computational complexity is comparable, TT-based option pricing holds an advantage in terms of computation time.
-This is because the TT-based method involves only multiplications and additions as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, iterated $dN_{\rm path}$ times.
-When the computational complexity of both methods was comparable, we confirmed that our proposed method was advantageous in terms of computation time.
-At this point, we also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths.
-However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\simeq 10$.
+Up to $d=11$, we demonstrated superiority in both computational complexity and time.
+Here, we argue that, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time.
+This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times.
+Indeed, in our demonstration, we confirmed that the TT-based method excels in computational complexity and is even more superior in terms of computational time.
+Note that both TT-based and MC-based option pricing implementations have room for improvement in computational time, and we consider the current comparison in time as reference data.
+We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths.
+However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\geq 12$.
 Now, let us consider how the proposed method provides benefits in practical business in financial firms.
 An expected way to utilize this method is as follows.
-At night when the financial market is closed, we learn the TTs, and then, during the day when the market is open, we use these TTs to quickly price the option for the fluctuating parameters.
+At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use these MPOs to quickly price the option for the fluctuating parameters.
 If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial.
-In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process.
+In light of this, it is reasonable that we compare the computational complexity in the TT-based method after the TTs are obtained with that in the MC-based method, neglecting the learning process and contraction.
 Finally, we discuss future research directions.
@@
 by increasing $N$ from the current value of 50 to, say, 100.
-It is a promising approach to incorporate parameter dependence using the Matrix Product Operators (MPO).
-It could eliminate the contraction between tensors related to parameters $p_{i}$ and $z_{i}$, potentially speeding up option pricing further.
+It is a promising approach to directly learning the MPOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing.
 At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself.
@@
 We are grateful to Marc K.
-Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the code for tensor cross interpolation.
+Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the tensor cross interpolation code.
+We deeply appreciate Marc K.
+Ritter for his important remarks on the further improvement of our proposed method.
 R.
@@
 thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor networks to option pricing.
+R.S.
+is grateful to Yusuke Himeoka, Yuta Mizuno, Wataru Mizukami, and Hiroshi Shinaoka, as the inspiration for this study was derived from their collaboration.
 K.

### v3

@@
 Unfortunately, running such a quantum algorithm requires fault-tolerant quantum computers, which may take decades to be developed.
-In light of this, applications of quantum-inspired algorithms to option pricing have been also studied, as solutions in the present or the near future [5, 6, 7, 8].
+In light of this, applications of quantum-inspired algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present or the near future [5, 6, 7, 8].
 Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6].
@@
 Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function.
-We then use these tensor networks to perform fast option pricing in response to various parameter changes.
+We use these tensor networks to perform fast option pricing in response to various parameter changes.
 To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices.
 In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice.
-When the computational complexity is smaller, our numerical results indicate that our method has a further advantage in terms of computation time, since our method involves only elementary arithmetic contrary to the Monte Carlo method involving more complicated operations.
+Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ for $d<10$ and by a factor of $O(10^{3})$ for about $d=10$ or $11$ (refer to FIG.
+2).
 We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths.
@@
 In this paper, we consider the underlying asset prices $\vec{S}(t)=(S_{1}(t),\cdots,S_{d}(t))$ in the Black-Scholes (BS) model described by the following stochastic differential equation
-$dS_{j}(t)=rS_{j}(t)dt+\sigma_{j}S_{j}(t)dW_{j}(t).$ (5)
+$dS_{m}(t)=rS_{m}(t)dt+\sigma_{m}S_{m}(t)dW_{m}(t).$ (5)
 Here, $W_{1}(t),\cdots,W_{d}(t)$ are the Brownian motions with constant correlation matrix $\rho_{ij}$, namely
-$dW_{i}(t)dW_{j}(t)=\rho_{ij}dt,$ (10)
+$dW_{m}(t)dW_{n}(t)=\rho_{mn}dt,$ (10)
 where $r\in\mathbb{R}$ and $\sigma_{1},\cdots,\sigma_{d}>0$ are constant parameters called the risk-free interest rate and the volatilities, respectively.
@@
 $q(\vec{x}|\vec{x}_{0})=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\exp\left(-\frac{1}{2}\left(\vec{x}-\vec{\mu}\right)^{T}\Sigma^{-1}\left(\vec{x}-\vec{\mu}\right)\right),$ (7)
-where $\Sigma:=(\sigma_{i}\sigma_{j}\rho_{ij}T)_{ij}$ is the covariance matrix of $\vec{x}$ and $\vec{\mu}:=\vec{x}_{0}+\left(rT-\frac{1}{2}\sigma_{1}^{2}T,\cdots,rT-\frac{1}{2}\sigma_{d}^{2}T\right)$.
+where $\Sigma:=(\sigma_{m}\sigma_{n}\rho_{mn}T)_{mn}$ is the covariance matrix of $\vec{x}$ and $\vec{\mu}:=\vec{x}_{0}+\left(rT-\frac{1}{2}\sigma_{1}^{2}T,\cdots,rT-\frac{1}{2}\sigma_{d}^{2}T\right)$.
 Note that, in Eq.
@@
 is the characteristic function, and in the BS model, it is given by
-$\phi(\vec{z})=\exp\left(i\sum_{j=1}^{d}z_{j}\mu_{j}-\frac{T}{2}\sum_{j=1}^{d}\sum_{k=1}^{d}\sigma_{j}\sigma_{k}z_{j}z_{k}\rho_{jk}\right).$ (10)
+$\phi(\vec{z})=\exp\left(i\sum_{m=1}^{d}z_{m}\mu_{m}-\frac{T}{2}\sum_{m=1}^{d}\sum_{k=1}^{d}\sigma_{m}\sigma_{k}z_{m}z_{k}\rho_{mk}\right).$ (10)
 […]
@@
 and its Fourier transform is [26]
-$\hat{v}_{\min}(\vec{z})=-\frac{K^{1+i\sum_{j=1}^{d}z_{j}}}{(-1)^{d}\left(1+i\sum_{j=1}^{d}z_{j}\right)\prod_{j=1}^{d}(iz_{j})}.$ (12)
+$\hat{v}_{\min}(\vec{z})=-\frac{K^{1+i\sum_{m=1}^{d}z_{m}}}{(-1)^{d}\left(1+i\sum_{m=1}^{d}z_{m}\right)\prod_{m=1}^{d}(iz_{m})}.$ (12)
 […]
@@
 $\vec{z}_{\mathrm{gr},\vec{j}}:=(\eta_{1}(\vec{p})j_{1},\ldots,\eta_{d}(\vec{p})j_{d}),$ (14)
-where $\eta_{j}$ is the integration step size in the $j$-th direction.
-$\eta_{j}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{j}(\vec{p})$ in (14).
+where $\eta_{m}$ is the integration step size in the $m$-th direction.
+$\eta_{m}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{m}(\vec{p})$ in (14).
 In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec.
 V for the detail.
-$\Delta_{\rm vol}:=\Pi_{j=1}^{d}\eta_{j}$ is the volume element.
+$\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
 III.2 Fourier transform-based option pricing with tensor trains
@@
 (13) by
-$V(\vec{\sigma},\vec{S}_{0})\simeq\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\tilde{\phi}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}}.$ (15)
+$V(\vec{\sigma},\vec{S}_{0})\simeq\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\tilde{\phi}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}}\Delta_{\mathrm{vol}}.$ (15)
 Thanks to TCI, we can obtain the approximate TTs avoiding the evaluations of $\phi$ and $\hat{v}$ at all the grid points.
-Besides, given the TTs, we can compute the sum in (15) as the contraction of two TTs without exponentially many iterations: with the bond dimensions at most $\chi$, the number of multiplications and additions is of order $O(d\chi^{3})$.
+Besides, given the TTs, we can compute the sum in (15) as the contraction of two TTs without exponentially many iterations: with the bond dimensions at most $\chi$, the number of multiplications and additions is of order $O(dN\chi^{3})$.
 Hereafter, we simply call this approach for FT-based option pricing aided by TTs TT-based option pricing.
@@
 In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
-IV Learning parameter dependence with tensor trains
+IV Learning parameter dependence with tensor networks
 IV.1 Outline
-Option prices depend on input parameters such as the volatilities $\vec{\sigma}$ and the present asset prices $\vec{S}_{0}$.
+Option prices $V(\vec{p})$ depend on input parameters $\vec{p}$ such as the volatilities $\vec{\sigma}$ and the present asset prices $\vec{S}_{0}$.
 In the rapidly changing financial market, these input parameters vary from time to time, which causes the change of the option price.
-Therefore, if we have a function that efficiently outputs an accurate approximation of the option price for various values of the input parameters, it provides a large benefit to practical business.
+Therefore, if we have a function, i.e., tensor networks, that efficiently outputs an accurate approximation of the option price for various values of the input parameters, it provides a large benefit to practical business.
 Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set.
@@
 1.
+Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$.
 Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$.
@@
 We reduce the bond dimensions of these TTs using SVD.
-Subsequently, we contract adjacent core tensors pairwise and obtain matrix product operators (MPOs).
-Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$.
-Having these MPOs, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig.
+We then contract adjacent core tensors pairwise and obtain two matrix product operators (MPOs) $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$.
+We contract these MPOs along the index vector $\vec{j}$ as
+$K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
+and optimize the bond dimensions of this resultant MPO $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ using SVD.
+Having this MPO, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig.
 1 (b).
-By fixing the index $\vec{k}$ of $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ to the value corresponding to the specified $\vec{p}$, we get the $d$-dimensional TTs $\tilde{\phi}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$.
-Contracting these TTs as Eq.
-(15), we get the option price for the specified $\vec{p}$.
+By fixing the index $\vec{k}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
+$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
 Here, we note that the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ for $\hat{v}$ and its TT approximation $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ corresponding to $\vec{p}$, which is now $\vec{\sigma}$ or $\vec{S}_{0}$, the parameter not affecting $\hat{v}$.
@@
 Thus, we will leave such a study for future work.
-Figure 1: Fast option pricing based on TTs proposed in this study.
+Figure 1: Fast option pricing based on TNs proposed in this study.
 In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD.
-Then, we contract the tensors associated with $p_{i}$ and $z_{i}$, resulting in MPOs.
-In (b), we use these MPOs to perform fast option pricing for a specified parameter $\vec{p}$.
+We then contract the tensors associated with $p_{m}$ and $z_{m}$, resulting in two MPOs.
+We then take the contraction of these two MPOs with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting MPO using SVD.
+In (b), we use this MPO to perform fast option pricing for a specified parameter $\vec{p}$.
 IV.2 Computational complexity
-The computational complexity of TT-based option pricing, involving computation of the inner product between two TTs, is $O(d\chi^{3}N)$.
-In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations.
+The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{K}^{2})$ [25].
+Here, we denote the maximum bond dimensions as $\chi_{K}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations.
 Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec.
 V.
-Here, we ignore the computational complexity of learning the TTs by TCI, SVD, and contraction of tensors associated with $p_{i}$ and $z_{i}$ and consider that after we get the MPOs only.
+Here, we ignore the computational complexity of all the processes in FIG.
+1 (a) and consider that in FIG.
+1 (b) after we get the MPO only.
 This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing.
@@
 In the following, we describe the parameter values used in this study, the software used, and how the errors were evaluated.
-Ranges of ${\sigma_{i}}$ and ${S^{i}_{0}}$
+Ranges of ${\sigma_{m}}$ and ${S_{m,0}}$
 With respect to $\vec{p}$, on which the TTs learn the dependence of the functions $\phi$ and $\hat{v}$, we take the two test cases: $\vec{p}$ is the volatilities $\vec{\sigma}$ or the present asset prices $\vec{S}_{0}$.
 In the proposed method, we need to set the range in the space of $\vec{p}$ and the grid points in it.
-For each volatility $\sigma_{i}$, we set the range to $\sigma_{i}\in[0.15,0.25]$, where the center $\sigma_{i}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [28] and the range width ±$\pm 0.05$ covers the changes in this index on most days.
-For each $S_{i,0}$, we set the range to $S_{i,0}\in[90,120]$, which corresponds to the 20% variation of the asset price centered at 100.
+For each volatility $\sigma_{m}$, we set the range to $\sigma_{m}\in[0.15,0.25]$, where the center $\sigma_{m}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [28] and the range width ±$\pm 0.05$ covers the changes in this index on most days.
+For each $S_{m,0}$, we set the range to $S_{m,0}\in[90,120]$, which corresponds to the 20% variation of the asset price centered at 100.
 The lower bound is set to not 80 but 90 because the price of the option we take as an example is negligibly small for $S_{0}<90$.
-For both $\sigma_{i}$ and $S_{i,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
+For both $\sigma_{m}$ and $S_{m,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
 Other parameters
 The other parameters for option pricing are fixed to the values summarized in the table below.
-$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{ii}$ $\rho_{ij(i\neq j)}$ $N$
+$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{mm}$ $\rho_{mn(m\neq n)}$ $N$
 $1$ $0.01$ $100$ $100$ $5/d$ $1$ $1/3$ $50$
@@
 The spread of the characteristic function changes depending on the values of the parameters $\vec{p}$ and the number of assets.
-Accordingly, it is necessary to adjust the grid interval $\eta_{j}$ for integration in the wavenumber space.
+Accordingly, it is necessary to adjust the grid interval $\eta_{m}$ for integration in the wavenumber space.
 For example, consider the case where $\vec{p}=\vec{\sigma}$.
 From Eq.
-(10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{j}z_{j}\lesssim 1$ for each $j$.
+(10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{m}z_{m}\lesssim 1$ for each $m$.
 It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval.
-Currently, as $z_{j}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{j}$, $\eta_{j}$ and $\sigma_{j}$ must satisfy $\sigma_{j}\eta_{j}\ll 1$ and $\sigma_{j}\times N\eta_{j}\gg 1$.
-Therefore, in general, we need the adjustment that $\eta_{j}$ increases when $\sigma$ decreases and vice versa.
-Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{j}$ set to some constant, and thus we did so.
-In the cases with $d=10,11$, $\eta_{j}$ is adjusted as
-$\eta_{j}=\frac{a}{\sigma_{j}}+b$ (17)
-To set the constants $a$ and $b$, we ran TT-based option pricing for several values of $\sigma_{j}$ adjusting $\eta_{j}$, found the pairs $(\sigma_{j},\eta_{j})$ that yields the accurate price, and adjust $a$ and $b$ so that Eq.
-(17) approximately reproduces such $\eta_{j}$ from $\sigma_{j}$.
-Also when $\vec{p}=\vec{S}_{0}$, we set $\eta_{j}$ to some constant in almost all the test cases. Only for $d=11$, we try the adjustment that
-$\eta_{j}=\frac{\alpha}{S_{j,0}}+\beta$ (18)
-with $\alpha$ and $\beta$ chosen according to the option pricing result for several $\eta_{j}$ and $\sigma_{j}$.
-In general, larger $S_{j,0}$, which means larger $x_{j,0}$, makes $\phi$ as a function of $z_{j}$ oscillate more rapidly, and thus we need finer grid points to resolve such an oscillation.
-Note that when $\eta_{j}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.
+Currently, as $z_{m}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{m}$, $\eta_{m}$ and $\sigma_{m}$ must satisfy $\sigma_{m}\eta_{m}\ll 1$ and $\sigma_{m}\times N\eta_{m}\gg 1$.
+Therefore, in general, we need the adjustment that $\eta_{m}$ increases when $\sigma$ decreases and vice versa.
+Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{m}$ set to some constant, and thus we did so.
+In the cases with $d=10,11$, $\eta_{m}$ is adjusted as
+$\eta_{m}=\frac{a}{\sigma_{m}}+b$ (19)
+To set the constants $a$ and $b$, we ran TT-based option pricing for several values of $\sigma_{m}$ adjusting $\eta_{m}$, found the pairs $(\sigma_{m},\eta_{m})$ that yields the accurate price, and adjust $a$ and $b$ so that Eq.
+(19) approximately reproduces such $\eta_{m}$ from $\sigma_{m}$.
+Also when $\vec{p}=\vec{S}_{0}$, we set $\eta_{m}$ to some constant in almost all the test cases. Only for $d=11$, we try the adjustment that
+$\eta_{m}=\frac{\alpha}{S_{m,0}}+\beta$ (20)
+with $\alpha$ and $\beta$ chosen according to the option pricing result for several $\eta_{m}$ and $\sigma_{m}$.
+In general, larger $S_{m,0}$, which means larger $x_{m,0}$, makes $\phi$ as a function of $z_{m}$ oscillate more rapidly, and thus we need finer grid points to resolve such an oscillation.
+Note that when $\eta_{m}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.
 Error evaluation
@@
 The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively.
-Also, the computational times of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively.
+Also, the computational time of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively.
 To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$.
-Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{v}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
-The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi^{\phi}_{\mathrm{max}}$ and $\chi^{\hat{v}}_{\mathrm{max}}$, respectively, in TABLE 2.
+Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{\hat{v}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
+The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2.
+We set the tolerance of SVD for obtaining $\chi_{K}$, maximum bond dimension of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{K}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
 V.2.1 The case of varying $\vec{\sigma}$
@@
 3(a).
-By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ could generally be maintained at less than 10 (refer to TABLE 2).
+By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at less than 10.
 The details of this compression by SVD are described in Appendix A.
-Furthermore, The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
-We argue that, when the computational complexity of the two approaches is at least comparable, the TT-based option pricing has an advantage in terms of computational time.
-This is because, in the MC-based method, the unit operations such as random number generation are more complex than those in the TT-based method, which are just multiplications and additions.
-Indeed, in our demonstration, we have shown that the TT-based method is advantageous in terms of computational complexity and exhibits even better performance in terms of computational time.
-However, it is important to note that the actual wall clock time may vary depending on the quality of the implementation.
+The maximum bond dimension $\chi_{K}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of about 5, when $\eta$ is constant, but increases to about 20 when $\eta$ is automatically determined for each asset, which leads to increased computational complexity at around $d=10$.
+Still, the computational complexity is much lower than that of MC-based option pricing.
+The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
 It has been observed that accuracy tends to decrease as $d$ increases.
@@
 (a) $\vec{\sigma}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
-$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $4.76\times 10^{4}$ $5.0\times 10^{5}$ $3.27\times 10^{-4}$ $2.84\times 10^{-2}$ $4$ $7$
-$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $7.94\times 10^{4}$ $6.0\times 10^{5}$ $6.11\times 10^{-4}$ $3.38\times 10^{-2}$ $5$ $7$
-$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0202 0.0236 $1.21\times 10^{5}$ $7.0\times 10^{5}$ $7.66\times 10^{-4}$ $3.71\times 10^{-2}$ $6$ $7$
-$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0180 0.0253 $1.48\times 10^{5}$ $8.0\times 10^{5}$ $9.07\times 10^{-4}$ $4.09\times 10^{-2}$ $6$ $7$
-$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $2.01\times 10^{5}$ $9.0\times 10^{5}$ $1.24\times 10^{-3}$ $4.51\times 10^{-2}$ $7$ $7$
-$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0308 0.0357 $2.48\times 10^{5}$ $1.0\times 10^{6}$ $2.19\times 10^{-3}$ $5.11\times 10^{-2}$ $7$ $7$
-$11$ auto $3.5\times 10^{-6}$ $3.5\times 10^{-6}$ 0.0341 0.0347 $3.02\times 10^{5}$ $1.1\times 10^{6}$ $2.57\times 10^{-3}$ $5.04\times 10^{-2}$ $8$ $7$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{K}$
+$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $3.12\times 10^{-5}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
+$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $3.58\times 10^{-5}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
+$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.13\times 10^{-5}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
+$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $4.58\times 10^{-5}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
+$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.15\times 10^{-5}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
+$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0309 0.0357 $2359$ $1.0\times 10^{6}$ $6.68\times 10^{-4}$ $5.11\times 10^{-2}$ $7$ $7$ $19$
+$11$ auto $3.5\times 10^{-6}$ $3.5\times 10^{-6}$ 0.0341 0.0347 $2797$ $1.1\times 10^{6}$ $7.91\times 10^{-4}$ $5.04\times 10^{-2}$ $8$ $7$ $20$
 (b) $\vec{S}_{0}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi^{\phi}_{\mathrm{}}$ $\chi^{v}_{\mathrm{}}$
-$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $3.53\times 10^{4}$ $5.0\times 10^{5}$ $2.47\times 10^{-4}$ $2.90\times 10^{-2}$ $4$ $6$
-$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $6.38\times 10^{4}$ $6.0\times 10^{5}$ $5.61\times 10^{-4}$ $3.28\times 10^{-2}$ $5$ $6$
-$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $1.00\times 10^{5}$ $7.0\times 10^{5}$ $6.80\times 10^{-4}$ $3.66\times 10^{-2}$ $6$ $7$
-$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0258 0.0340 $1.54\times 10^{5}$ $8.0\times 10^{5}$ $9.71\times 10^{-4}$ $3.93\times 10^{-2}$ $7$ $7$
-$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $1.99\times 10^{5}$ $9.0\times 10^{5}$ $1.37\times 10^{-3}$ $4.54\times 10^{-2}$ $7$ $7$
-$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $3.00\times 10^{5}$ $1.0\times 10^{6}$ $1.88\times 10^{-3}$ $4.93\times 10^{-2}$ $7$ $8$
-$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0377 0.0401 $7.33\times 10^{5}$ $1.1\times 10^{6}$ $6.25\times 10^{-3}$ $5.28\times 10^{-2}$ $9$ $14$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{K}$
+$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.37\times 10^{-5}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
+$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.26\times 10^{-5}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
+$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $3.61\times 10^{-5}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
+$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.25\times 10^{-5}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
+$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $5.12\times 10^{-5}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
+$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $6.02\times 10^{-5}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
+$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0379 0.0401 $853$ $1.1\times 10^{6}$ $3.43\times 10^{-4}$ $5.28\times 10^{-2}$ $9$ $14$ $11$
 Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
-Here, we set the ranges $\sigma_{i}\in[0.15,0.25]$ and $S_{0}^{i}\in[90,120]$, and place 100 equally spaced grid points within these ranges.
+Here, we set the ranges $\sigma_{m}\in[0.15,0.25]$ and $S_{m,0}\in[90,120]$, and place 100 equally spaced grid points within these ranges.
 The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$.
-The term auto in (a) and (b) indicate that $\eta_{i}$ was automatically determined for each asset based on Eqs.
+The term auto in (a) and (b) indicate that $\eta_{m}$ was automatically determined for each asset based on Eqs.
 (19) and (20), respectively.
+The computational complexity and time increases when $\eta$ is set by Eq.
+(19) or (20).
 $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
@@
 We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs.
-In TT-based option pricing with $\vec{S_{0}}$, the sharp increase in computational complexity at $d=11$ arises.
-This stems from the larger bond dimensions with a sufficiently low cutoff value for SVD in order to maintain its accuracy
+In TT-based option pricing with $\vec{\sigma}$ and $\vec{S_{0}}$, the sharp increase in computational complexity at $d=10$ and $d=11$ arises, respectively.
+This stems from the necessity to set $\eta$ by Eq.
+(19) or (20), which in turn makes the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{m}$ for $\vec{p}$.
 Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$.
-The odd bond $l=2i-1$ connects the core tensors for $j_{i}$ and $k_{i}$, the indexes for $z_{i}$ and $p_{i}$, respectively, and the even bond $l=2i$ connects those for $k_{i}$ and $j_{i+1}$.
-In the cases corresponding to the left panels and the bottom right panel, $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{i}$ and $z_{i+1}$).
+The odd bond $l=2i-1$ connects the core tensors for $j_{m}$ and $k_{m}$, the indexes for $z_{m}$ and $p_{m}$, respectively, and the even bond $l=2i$ connects those for $k_{m}$ and $j_{m+1}$.
+In the cases corresponding to the left panels and the bottom right panel, $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{m}$ and $z_{m+1}$).
 It is noteworthy that the graph of bond dimensions exhibits a characteristic jagged shape.
-The bond dimensions between sites for $j_{i}$ and $k_{i}$, which are related to the same asset, are large, and the ones between sites for $k_{i}$ and $j_{i+1}$, which are related to different assets, are small.
+The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
 V.2.2 The case of varying $\vec{S_{0}}$
@@
 TT-based option pricing demonstrates superiority over the Monte Carlo method in terms of computational complexity and time.
-For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{i}$ and $z_{i+1}$ were reduced to around 10 (refer to Fig.
+For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ were reduced to around 10 (refer to Fig.
 3(b)).
-However, at $d$ = 11, the computational complexity of TT-based calculation increases significantly.
-This is because it was necessary to increase the bond dimension by setting the SVD cutoff sufficiently low to maintain accuracy (refer to TABLE 2).
-This suggests that $d=11$ may already be reaching the upper limit of $d$ where accuracy can be maintained while surpassing the computational complexity over MC-based calculations.
+The maximum bond dimension, $\chi_{K}$, of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is about 6 when $\eta$ is constant, while it increases to 11 when $\eta$ is adjusted per asset, leading to higher computational complexity at about $d=11$.
+Nonetheless, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
 Compared to varying $\sigma$, its accuracy is worse at $d=10$.
@@
 V.2.4 Total computational time for obtaining the MPOs
-Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ through TCI and SVD.
+Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ and $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
 We focus on TCI because it dominates over SVD.
@@
 Up to $d=11$, we demonstrated superiority in both computational complexity and time.
-Here, we argue that, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time.
+Here, we argue that, in principle, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time.
 This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times.
-Indeed, in our demonstration, we confirmed that the TT-based method excels in computational complexity and is even more superior in terms of computational time.
-Note that both TT-based and MC-based option pricing implementations have room for improvement in computational time, and we consider the current comparison in time as reference data.
+However, in our demonstration, compared to the huge computational complexity advantage, the computational time advantage of TT-based option pricing is milder.
+This is because the TT-based method is bottlenecked by specifying the values of $\vec{p}$.
+Specifically, accessing the memory for the specific tensor components $K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ dominates the computational time for additions and multiplications in tensor contraction.
+Therefore, in the TT-based method, and perhaps the MC-based method too, the implementation may have room for improvement to reduce computational time.
 We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths.
@@
 An expected way to utilize this method is as follows.
-At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use these MPOs to quickly price the option for the fluctuating parameters.
+At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the MPO to quickly price the option for the fluctuating parameters.
 If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial.
@@
 At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself.
-For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TT format is desired, although this study tried to incorporate either of them as the first step.
+For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step.
 Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.
@@
 We deeply appreciate Marc K.
-Ritter for his important remarks on the further improvement of our proposed method.
+Ritter and Hiroshi Shinaoka for their critical remarks on the further improvement of our proposed method.
 R.
@@
 The left vertical axis represents $r_{\mathrm{TT}}$, while the right vertical axis shows $c_{\mathrm{TT}}$.
-Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
-The cutoff $\epsilon_{\text{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$.
+Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon^{\phi}_{\mathrm{SVD}}$(=$\epsilon^{\hat{v}}_{\mathrm{SVD}}$) to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
+Here, we fixed the tolerance of SVD $\epsilon^{K}_{\mathrm{SVD}}$ to $10^{-10}$.
+The cutoff $\epsilon^{\phi}_{\mathrm{SVD}}$(=$\epsilon^{\hat{v}}_{\mathrm{SVD}}$) for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$.
 From Fig.
 4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ while minimizing the computational complexity.
+By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
 From the fact that we can keep the maximum relative error $r_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information.

### v4

@@
 We show that, in the tested cases involving up to 11 assets, the proposed method outperforms MC-based option pricing with $10^{5}$ paths in terms of computational complexity while keeping comparable accuracy.
-I Introduction
+1 Introduction
 Financial firms are conducting demanding numerical calculations in their business.
@@
 On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews.
-and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17], and differential equations [7, 18].
+and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17, 18], and partial differential equations [7, 19].
 Ref.
@@
 According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
-To address these issues, we focus on another usage of tensor networks that can embed parameter dependence [22, 23, 24] to make FT-based option pricing more efficient.
+To address this issue, we focus on another usage of tensor networks that can embed parameter dependence [23, 24, 25] to make FT-based option pricing more efficient.
 Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function.
@@
 VI.
-II Tensor train
+2 Tensor train
 A $d$-way tensor $F_{x_{1},\dots,x_{d}}$, where each local index $x_{l}$, $l=1,\dots,d$, has a local dimension $N$, can be decomposed into a TT format with a low-rank structure.
@@
 This is an equivalent expression to the wave function $F_{x_{1},\ldots,x_{d}}$ of a quantum system with $d$ $N$-level qudits as follows:
-|$\ket{F_{x_{1},x_{2},\ldots,x_{d}}}$
-$=\sum_{l_{1}}^{\chi_{1}}\cdots\sum_{l_{d-1}}^{\chi_{d}}F_{l_{1},x_{1}}^{(1)}F_{l_{1}l_{2},x_{2}}^{(2)}\cdots F_{l_{d-1},x_{d}}^{(d)}\ket{\vec{x}},$ (2)
+|$\ket{F_{x_{1},x_{2},\ldots,x_{d}}}=\sum_{l_{1}}^{\chi_{1}}\cdots\sum_{l_{d-1}}^{\chi_{d}}F_{l_{1},x_{1}}^{(1)}F_{l_{1}l_{2},x_{2}}^{(2)}\cdots F_{l_{d-1},x_{d}}^{(d)}\ket{\vec{x}},$ (2)
 […]
-II.1 Compression tenchniques
+2.1 Compression tenchniques
 We introduce the two compression techniques used in this study.
-II.1.1 Tensor cross interpolation
+2.1.1 Tensor cross interpolation
 Tensor cross interpolation (TCI) is a technique to compress tensors corresponding to discretized multivariate functions with a low-rank TT representation.
@@
 In this study, when we learn TT from functions with TCI, we add the pivots so that the error in the maximum norm $(\epsilon_{\mathrm{TCI}})$ is to be minimized.
-$\epsilon_{\mathrm{max}}=\frac{\|F_{\mathrm{TT}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{max}}}{\|F_{\mathrm{TT}}\|_{\mathrm{max}}}$ (3)
-where $F_{\mathrm{TT}}$ is a target tensor, $\tilde{F}_{\mathrm{TT}}$ is a low-rank approximation, and the maximum norm is evaluated as the maximum of the absolute values of the entries at the pivots selected already.
+$\epsilon_{\mathrm{max}}=\frac{\|F_{x_{1},x_{2},\ldots,x_{d}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{max}}}{\|F_{x_{1},x_{2},\ldots,x_{d}}\|_{\mathrm{max}}}$ (3)
+where $F_{x_{1},x_{2},\ldots,x_{d}}$ is a target tensor [^5] Here, we assume that we can access the arbitrary elements of the target tensor.
+Note that we do not need to store all the elements of the target tensor., $\tilde{F}_{\mathrm{TT}}$ is a low-rank approximation, and the maximum norm is evaluated as the maximum of the absolute values of the entries at the pivots selected already.
 The computational complexity of TCI is roughly proportional to the number of elements in the TT, which is $O(d\chi^{2}N)$ with $\chi_{1},\cdots,\chi_{d}$ fixed to $\chi$.
-In addition, considering the case that zero is included in the reference function value, the error should be normalized by $\|F_{\mathrm{TT}}\|_{\mathrm{max}}$.
-II.1.2 Singular value decomposition
+In addition, considering the case that zero is included in the reference function value, the error should be normalized by $\|F_{x_{1},x_{2},\ldots,x_{d}}\|_{\mathrm{max}}$.
+2.1.2 Singular value decomposition
 In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low.
-The original TT, denoted as $F_{\mathrm{TT}}$, can be approximated by another TT, $\tilde{F}_{\mathrm{TT}}$, with reduced (optimized) bond dimensions such that
+The TT obtained from TCI, denoted as $F_{\mathrm{TT}}$, can be approximated by another TT, $\tilde{F}_{\mathrm{TT}}$, with reduced (optimized) bond dimensions such that
 $\frac{\|F_{\mathrm{TT}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{F}}^{2}}{\|F_{\mathrm{TT}}\|_{\mathrm{F}}^{2}}<\epsilon_{\mathrm{SVD}},$ (4)
@@
 [26].
-III Fourier transform-based option pricing aided by tensor cross interpolation
-III.1 Fourier transform-based option pricing
+3 Fourier transform-based option pricing aided by tensor cross interpolation
+3.1 Fourier transform-based option pricing
 In this paper, we consider the underlying asset prices $\vec{S}(t)=(S_{1}(t),\cdots,S_{d}(t))$ in the Black-Scholes (BS) model described by the following stochastic differential equation
@@
 In the numerical calculation of Eq. (8), we approximate it by discretization:
-$V(\vec{p})$
-$=\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\phi(-\vec{z}_{\mathrm{gr},\vec{j}}-i\vec{\alpha})\hat{v}(\vec{z}_{\mathrm{gr},\vec{j}}+i\vec{\alpha})\Delta_{\rm vol}.$ (13)
+$V(\vec{p})=\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\phi(-\vec{z}_{\mathrm{gr},\vec{j}}-i\vec{\alpha})\hat{v}(\vec{z}_{\mathrm{gr},\vec{j}}+i\vec{\alpha})\Delta_{\rm vol}.$ (13)
 Here, the even natural number $N$ is the number of the grid points in one dimension. $\vec{z}_{\mathrm{gr},\vec{j}}$ is the grid point specified by the integer vector $\vec{j}$ as
@@
 In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec.
-V for the detail.
+5 for the detail.
 $\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
-III.2 Fourier transform-based option pricing with tensor trains
+3.2 Fourier transform-based option pricing with tensor trains
 Note that to compute the sum in Eq.
@@
 Hereafter, we simply call this approach for FT-based option pricing aided by TTs TT-based option pricing.
-III.3 Monte Carlo-based option pricing
+3.3 Monte Carlo-based option pricing
 Here, we also make a brief description of the Monte Carlo (MC)-based option pricing.
@@
 In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
-IV Learning parameter dependence with tensor networks
-IV.1 Outline
+4 Learning parameter dependence with tensor networks
+4.1 Outline
 Option prices $V(\vec{p})$ depend on input parameters $\vec{p}$ such as the volatilities $\vec{\sigma}$ and the present asset prices $\vec{S}_{0}$.
@@
 We contract these MPOs along the index vector $\vec{j}$ as
-$K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
-and optimize the bond dimensions of this resultant MPO $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ using SVD.
+$O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
+and optimize the bond dimensions of this resultant MPO $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ using SVD.
 Having this MPO, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig.
 1 (b).
-By fixing the index $\vec{k}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
-$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
+By fixing the index $\vec{k}$ of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
+$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
 Here, we note that the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ for $\hat{v}$ and its TT approximation $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ corresponding to $\vec{p}$, which is now $\vec{\sigma}$ or $\vec{S}_{0}$, the parameter not affecting $\hat{v}$.
@@
 In the demonstration in Sec.
-V, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing.
+5, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing.
 On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains.
@@
 In the two test cases for our proposed method in Sec.
-V, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices.
+5, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices.
 Note that it is possible to include dependencies on other parameters in the TTs.
@@
 In (b), we use this MPO to perform fast option pricing for a specified parameter $\vec{p}$.
-IV.2 Computational complexity
-The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{K}^{2})$ [25].
+4.2 Computational complexity
+The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{O}^{2})$ [26].
 Here, we denote the maximum bond dimensions as $\chi_{K}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations.
@@
 As discussed in Sec.
-VI, we can reasonably find such a situation in practice.
-V Numerical demonstration
-V.1 Details
+6, we can reasonably find such a situation in practice.
+5 Numerical demonstration
+5.1 Details
 Now, as a demonstration, we apply the proposed method to pricing a $d$-asset European min-call option in the BS model.
@@
 Software and hardware used in this study
-The TCI code has not yet been made publicly available.
+TensorCrossInterpolation.jl for TCI [30] has not yet been made publicly available.
 Marc K.
-Ritter, Hiroshi Shinaoka, and Jan von Delft provided us with access to their TCI code.
+Ritter, Hiroshi Shinaoka, and Jan von Delft provided us with access to this.
 ITensors.jl [31] was utilized for computing the contraction and inner product of the tensor trains.
@@
 The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
-V.2 Results
+5.2 Results
 We show the results for the computational complexity, time, and accuracy of TT-based and MC-based option pricing when two parameters $\vec{\sigma}$ and $\vec{S_{0}}$ are varied.
@@
 To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$.
-Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{\hat{v}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
+Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
 The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2.
-We set the tolerance of SVD for obtaining $\chi_{K}$, maximum bond dimension of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{K}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
-V.2.1 The case of varying $\vec{\sigma}$
+We set the tolerance of SVD for obtaining $\chi_{K}$, maximum bond dimension of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{O}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
+5.2.1 The case of varying $\vec{\sigma}$
 TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence.
@@
 The details of this compression by SVD are described in Appendix A.
-The maximum bond dimension $\chi_{K}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of about 5, when $\eta$ is constant, but increases to about 20 when $\eta$ is automatically determined for each asset, which leads to increased computational complexity at around $d=10$.
+The maximum bond dimension $\chi_{O}$ of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of about 5, when $\eta$ is constant, but increases to about 20 when $\eta$ is automatically determined for each asset, which leads to increased computational complexity at around $d=10$.
 Still, the computational complexity is much lower than that of MC-based option pricing.
@@
 (a) $\vec{\sigma}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{K}$
-$5$ 0.9 $6.0\times 10^{-5}$ $1.0\times 10^{-4}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $3.12\times 10^{-5}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
-$6$ 0.75 $5.0\times 10^{-5}$ $5.0\times 10^{-4}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $3.58\times 10^{-5}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
-$7$ 0.69 $1.8\times 10^{-5}$ $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.13\times 10^{-5}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
-$8$ 0.66 $2.0\times 10^{-5}$ $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $4.58\times 10^{-5}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
-$9$ 0.61 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.15\times 10^{-5}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
-$10$ auto $5.0\times 10^{-6}$ $5.0\times 10^{-6}$ 0.0309 0.0357 $2359$ $1.0\times 10^{6}$ $6.68\times 10^{-4}$ $5.11\times 10^{-2}$ $7$ $7$ $19$
-$11$ auto $3.5\times 10^{-6}$ $3.5\times 10^{-6}$ 0.0341 0.0347 $2797$ $1.1\times 10^{6}$ $7.91\times 10^{-4}$ $5.04\times 10^{-2}$ $8$ $7$ $20$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{O}$
+$5$ 0.9 $6.0\times 10^{-5}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $3.12\times 10^{-5}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
+$6$ 0.75 $5.0\times 10^{-5}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $3.58\times 10^{-5}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
+$7$ 0.69 $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.13\times 10^{-5}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
+$8$ 0.66 $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $4.58\times 10^{-5}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
+$9$ 0.61 $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.15\times 10^{-5}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
+$10$ auto $5.0\times 10^{-6}$ 0.0309 0.0357 $2359$ $1.0\times 10^{6}$ $6.68\times 10^{-4}$ $5.11\times 10^{-2}$ $7$ $7$ $19$
+$11$ auto $3.5\times 10^{-6}$ 0.0341 0.0347 $2797$ $1.1\times 10^{6}$ $7.91\times 10^{-4}$ $5.04\times 10^{-2}$ $8$ $7$ $20$
 (b) $\vec{S}_{0}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}^{\phi}$ $\epsilon_{\mathrm{SVD}}^{v}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{K}$
-$5$ 0.9 $2.0\times 10^{-4}$ $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.37\times 10^{-5}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
-$6$ 0.75 $7.0\times 10^{-5}$ $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.26\times 10^{-5}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
-$7$ 0.64 $3.0\times 10^{-5}$ $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $3.61\times 10^{-5}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
-$8$ 0.57 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.25\times 10^{-5}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
-$9$ 0.53 $1.0\times 10^{-5}$ $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $5.12\times 10^{-5}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
-$10$ 0.5 $3.0\times 10^{-6}$ $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $6.02\times 10^{-5}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
-$11$ auto $1.0\times 10^{-7}$ $1.0\times 10^{-7}$ 0.0379 0.0401 $853$ $1.1\times 10^{6}$ $3.43\times 10^{-4}$ $5.28\times 10^{-2}$ $9$ $14$ $11$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{O}$
+$5$ 0.9 $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.37\times 10^{-5}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
+$6$ 0.75 $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.26\times 10^{-5}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
+$7$ 0.64 $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $3.61\times 10^{-5}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
+$8$ 0.57 $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.25\times 10^{-5}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
+$9$ 0.53 $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $5.12\times 10^{-5}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
+$10$ 0.5 $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $6.02\times 10^{-5}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
+$11$ auto $1.0\times 10^{-7}$ 0.0379 0.0401 $853$ $1.1\times 10^{6}$ $3.43\times 10^{-4}$ $5.28\times 10^{-2}$ $9$ $14$ $11$
 Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
@@
 The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
-V.2.2 The case of varying $\vec{S_{0}}$
+5.2.2 The case of varying $\vec{S_{0}}$
 TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence.
@@
 3(b)).
-The maximum bond dimension, $\chi_{K}$, of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is about 6 when $\eta$ is constant, while it increases to 11 when $\eta$ is adjusted per asset, leading to higher computational complexity at about $d=11$.
+The maximum bond dimension, $\chi_{O}$, of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is about 6 when $\eta$ is constant, while it increases to 11 when $\eta$ is adjusted per asset, leading to higher computational complexity at about $d=11$.
 Nonetheless, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
 Compared to varying $\sigma$, its accuracy is worse at $d=10$.
-V.2.3 Randomness in learning the TTs
+5.2.3 Randomness in learning the TTs
 Here, we mention the randomness in learning the TTs and the error induced by it.
@@
 The aforementioned mean and standard deviation are computed from successful cases.
-V.2.4 Total computational time for obtaining the MPOs
-Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ and $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
+5.2.4 Total computational time for obtaining the MPOs
+Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ and $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
 We focus on TCI because it dominates over SVD.
@@
 This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec.
-VI.
+6.
 We also note that, when we do not incorporate the parameter dependence into the tensor networks for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence.
@@
 This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
-VI Summary and discussion
+6 Summary and discussion
 We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs.
@@
 This is because the TT-based method is bottlenecked by specifying the values of $\vec{p}$.
-Specifically, accessing the memory for the specific tensor components $K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ dominates the computational time for additions and multiplications in tensor contraction.
+Specifically, accessing the memory for the specific tensor components $O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ dominates the computational time for additions and multiplications in tensor contraction.
 Therefore, in the TT-based method, and perhaps the MC-based method too, the implementation may have room for improvement to reduce computational time.
@@
 Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.
-Acknowledgements.
+Acknowledgements
 We are grateful to Marc K.
-Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to the tensor cross interpolation code.
+Ritter, Hiroshi Shinaoka, and Jan von Delft for providing access to TensorCrossInterpolation.jl for TCI [30].
 We deeply appreciate Marc K.
@@
 R.S.
-is grateful to Yusuke Himeoka, Yuta Mizuno, Wataru Mizukami, and Hiroshi Shinaoka, as the inspiration for this study was derived from their collaboration.
+is grateful to Yusuke Himeoka, Yuta Mizuno, Wataru Mizukami, and Hiroshi Shinaoka since R.S.
+got the inspiration for this study through their collaboration.
 K.
@@
 JPMJPF2014.
-Appendix A How to set the tolerance of SVD
+Appendix A How to set the tolerance $\epsilon_{\mathrm{SVD}}$
 Figure 4: The maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with various $\epsilon_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
 The left vertical axis represents $r_{\mathrm{TT}}$, while the right vertical axis shows $c_{\mathrm{TT}}$.
-Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon^{\phi}_{\mathrm{SVD}}$(=$\epsilon^{\hat{v}}_{\mathrm{SVD}}$) to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
-Here, we fixed the tolerance of SVD $\epsilon^{K}_{\mathrm{SVD}}$ to $10^{-10}$.
-The cutoff $\epsilon^{\phi}_{\mathrm{SVD}}$(=$\epsilon^{\hat{v}}_{\mathrm{SVD}}$) for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$.
+Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
+Here, we fixed the tolerance of SVD $\epsilon^{O}_{\mathrm{SVD}}$ to $10^{-10}$.
+The tolerance $\epsilon_{\mathrm{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$.
 From Fig.

### v5

@@
-Learning parameter dependence for Fourier-based option pricing with tensor networks
+Learning parameter dependence for Fourier-based option pricing with tensor trains
 Abstract
 A long-standing issue in mathematical finance is the speed-up of option pricing, especially for multi-asset options.
-A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor networks to compress high-dimensional tensors.
-Another usage of the tensor network is to compress functions, including their parameter dependence.
-Here, we propose a pricing method, where, by a tensor learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters.
+A recent study has proposed to use tensor train learning algorithms to speed up Fourier transform (FT)-based option pricing, utilizing the ability of tensor trains to compress high-dimensional tensors.
+Another usage of the tensor train is to compress functions, including their parameter dependence.
+Here, we propose a pricing method, where, by a tensor train learning algorithm, we build tensor trains that approximate functions appearing in FT-based option pricing with their parameter dependence and efficiently calculate the option price for the varying input parameters.
 As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices.
-We show that, in the tested cases involving up to 11 assets, the proposed method outperforms MC-based option pricing with $10^{5}$ paths in terms of computational complexity while keeping comparable accuracy.
-1 Introduction
+We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ paths in terms of computational complexity while keeping comparable accuracy.
+I Introduction
 Financial firms are conducting demanding numerical calculations in their business.
@@
 In light of this, applications of quantum-inspired algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present or the near future [5, 6, 7, 8].
-Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6].
+Among them, this paper focuses on the application of tensor train (TT) [9]  [^3] Tensor train is a mathematically equivalent to matrix product state (MPS) [10].
+learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in [6].
 This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$.
@@
 On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^4] See [13, 14] as reviews.
-and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17, 18], and partial differential equations [7, 19].
+and recently, it has also been utilized in various fields such as machine learning [15, 16, 17], quantum field theory [18, 19, 20], and partial differential equations [7, 21, 22, 23, 24].
 Ref.
@@
 According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
-To address this issue, we focus on another usage of tensor networks that can embed parameter dependence [23, 24, 25] to make FT-based option pricing more efficient.
+To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [28, 29, 30] to make FT-based option pricing more efficient.
 Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function.
-We use these tensor networks to perform fast option pricing in response to various parameter changes.
+We use these tensor trains to perform fast option pricing in response to various parameter changes.
 To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices.
 In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice.
-Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ for $d<10$ and by a factor of $O(10^{3})$ for about $d=10$ or $11$ (refer to FIG.
+Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ up to $d=11$ (refer to FIG.
 2).
@@
 III, we review the FT-based option pricing and its reformulation aided by TCI.
-We propose a new scheme for fast option pricing by learning tensor networks with parameter dependence in Sec.
+We propose a new scheme for fast option pricing by learning tensor trains with parameter dependence in Sec.
 IV.
@@
 VI.
-2 Tensor train
+II Tensor train
 A $d$-way tensor $F_{x_{1},\dots,x_{d}}$, where each local index $x_{l}$, $l=1,\dots,d$, has a local dimension $N$, can be decomposed into a TT format with a low-rank structure.
@@
 […]
-2.1 Compression tenchniques
+II.1 Compression tenchniques
 We introduce the two compression techniques used in this study.
-2.1.1 Tensor cross interpolation
+II.1.1 Tensor cross interpolation
 Tensor cross interpolation (TCI) is a technique to compress tensors corresponding to discretized multivariate functions with a low-rank TT representation.
@@
 In addition, considering the case that zero is included in the reference function value, the error should be normalized by $\|F_{x_{1},x_{2},\ldots,x_{d}}\|_{\mathrm{max}}$.
-2.1.2 Singular value decomposition
+II.1.2 Singular value decomposition
 In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low.
-The TT obtained from TCI, denoted as $F_{\mathrm{TT}}$, can be approximated by another TT, $\tilde{F}_{\mathrm{TT}}$, with reduced (optimized) bond dimensions such that
-$\frac{\|F_{\mathrm{TT}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{F}}^{2}}{\|F_{\mathrm{TT}}\|_{\mathrm{F}}^{2}}<\epsilon_{\mathrm{SVD}},$ (4)
-where $\|\cdots\|_{\mathrm{F}}$ denotes the Frobenius norm.
-Note that the approximation error is conventionally expressed as the squared deviation.
-Given a tolerance $\epsilon_{\mathrm{SVD}}$, truncation via SVD results in an optimally low-rank TT approximation characterized by the smallest possible rank.
+This is done by first canonicalizing the TT using QR decompositions, and then performing the compression via SVD, discarding singular values that are smaller than the threshold $\epsilon_{\mathrm{SVD}}$ set for each bond.
 For more technical details, readers are referred to Ref.
-[26].
-3 Fourier transform-based option pricing aided by tensor cross interpolation
-3.1 Fourier transform-based option pricing
+[10, 9].
+III Fourier transform-based option pricing aided by tensor cross interpolation
+III.1 Fourier transform-based option pricing
 In this paper, we consider the underlying asset prices $\vec{S}(t)=(S_{1}(t),\cdots,S_{d}(t))$ in the Black-Scholes (BS) model described by the following stochastic differential equation
@@
 In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec.
-5 for the detail.
+V for the detail.
 $\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
-3.2 Fourier transform-based option pricing with tensor trains
+III.2 Fourier transform-based option pricing with tensor trains
 Note that to compute the sum in Eq.
@@
 Hereafter, we simply call this approach for FT-based option pricing aided by TTs TT-based option pricing.
-3.3 Monte Carlo-based option pricing
+III.3 Monte Carlo-based option pricing
 Here, we also make a brief description of the Monte Carlo (MC)-based option pricing.
@@
 In the MC simulation for $d$ assets with $N_{\rm path}$, the number of such operations is $O(dN_{\rm path})$, and we hereafter estimate the computational complexity of MC-based option pricing by this.
-4 Learning parameter dependence with tensor networks
-4.1 Outline
+IV Learning parameter dependence with tensor trains
+IV.1 Outline
 Option prices $V(\vec{p})$ depend on input parameters $\vec{p}$ such as the volatilities $\vec{\sigma}$ and the present asset prices $\vec{S}_{0}$.
 In the rapidly changing financial market, these input parameters vary from time to time, which causes the change of the option price.
-Therefore, if we have a function, i.e., tensor networks, that efficiently outputs an accurate approximation of the option price for various values of the input parameters, it provides a large benefit to practical business.
+Therefore, if we have a function, i.e., tensor trains, that efficiently outputs an accurate approximation of the option price for various values of the input parameters, it provides a large benefit to practical business.
 Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set.
@@
 We reduce the bond dimensions of these TTs using SVD.
-We then contract adjacent core tensors pairwise and obtain two matrix product operators (MPOs) $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$.
-We contract these MPOs along the index vector $\vec{j}$ as
-$O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
-and optimize the bond dimensions of this resultant MPO $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ using SVD.
-Having this MPO, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig.
+We then contract adjacent core tensors pairwise and obtain two tensor train operators (TTOs), also known as matrix product operator (MPO), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$.
+We contract these TTOs along the index vector $\vec{j}$ as follows:
+$\tilde{V}^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (16)
+Note that this resultant TTO has only the diagonal elements with respect to the local indices $\vec{k}$.
+That is, when $k_{1}\neq k^{\prime}_{1},\ldots,k_{d}\neq k^{\prime}_{d}$, we set $\tilde{V}^{k_{1},\ldots,k_{d}}_{k^{\prime}_{1},\ldots,k^{\prime}_{d}}=0$.
+With this operation, the resultant tensor can be expressed as TT $\tilde{V}_{k_{1},\ldots,k_{d}}$.
+Then we optimize the bond dimensions of $\tilde{V}_{k_{1},\ldots,k_{d}}$ using SVD.
+Having this TT, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig.
 1 (b).
-By fixing the index $\vec{k}$ of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
-$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
+By fixing the index $\vec{k}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
+$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}\tilde{V}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (17)
 Here, we note that the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ for $\hat{v}$ and its TT approximation $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ corresponding to $\vec{p}$, which is now $\vec{\sigma}$ or $\vec{S}_{0}$, the parameter not affecting $\hat{v}$.
@@
 In the demonstration in Sec.
-5, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing.
+V, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing.
 On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains.
@@
 In the two test cases for our proposed method in Sec.
-5, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices.
+V, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices.
 Note that it is possible to include dependencies on other parameters in the TTs.
@@
 In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD.
-We then contract the tensors associated with $p_{m}$ and $z_{m}$, resulting in two MPOs.
-We then take the contraction of these two MPOs with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting MPO using SVD.
-In (b), we use this MPO to perform fast option pricing for a specified parameter $\vec{p}$.
-4.2 Computational complexity
-The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{O}^{2})$ [26].
-Here, we denote the maximum bond dimensions as $\chi_{K}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations.
-Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec.
-V.
+We then contract the tensors associated with $p_{m}$ and $z_{m}$, resulting in two TTOs.
+We then take the contraction of these two TTOs with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting TT using SVD.
+In (b), we use this TT to perform fast option pricing for a specified parameter $\vec{p}$.
+IV.2 Computational complexity
+[…]
 Here, we ignore the computational complexity of all the processes in FIG.
 1 (a) and consider that in FIG.
-1 (b) after we get the MPO only.
+1 (b) after we get the TT only.
 This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing.
 As discussed in Sec.
-6, we can reasonably find such a situation in practice.
-5 Numerical demonstration
-5.1 Details
+VI, we can reasonably find such a situation in practice.
+V Numerical demonstration
+V.1 Details
 Now, as a demonstration, we apply the proposed method to pricing a $d$-asset European min-call option in the BS model.
@@
 The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
-5.2 Results
+V.2 Results
 We show the results for the computational complexity, time, and accuracy of TT-based and MC-based option pricing when two parameters $\vec{\sigma}$ and $\vec{S_{0}}$ are varied.
@@
 Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
-The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2.
-We set the tolerance of SVD for obtaining $\chi_{K}$, maximum bond dimension of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{O}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
-5.2.1 The case of varying $\vec{\sigma}$
+The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2.
+We set the tolerance of SVD for obtaining $\chi_{V}$, maximum bond dimension of $\tilde{V}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
+V.2.1 The case of varying $\vec{\sigma}$
 TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence.
@@
 The details of this compression by SVD are described in Appendix A.
-The maximum bond dimension $\chi_{O}$ of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of about 5, when $\eta$ is constant, but increases to about 20 when $\eta$ is automatically determined for each asset, which leads to increased computational complexity at around $d=10$.
-Still, the computational complexity is much lower than that of MC-based option pricing.
+The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of 6 or less, regardless of when $\eta$ is constant or $\eta$ is automatically determined for each asset.
+Therefore, the computational complexity is much lower than that of MC-based option pricing.
 The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
@@
 (a) $\vec{\sigma}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{O}$
-$5$ 0.9 $6.0\times 10^{-5}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $3.12\times 10^{-5}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
-$6$ 0.75 $5.0\times 10^{-5}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $3.58\times 10^{-5}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
-$7$ 0.69 $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.13\times 10^{-5}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
-$8$ 0.66 $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $4.58\times 10^{-5}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
-$9$ 0.61 $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.15\times 10^{-5}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
-$10$ auto $5.0\times 10^{-6}$ 0.0309 0.0357 $2359$ $1.0\times 10^{6}$ $6.68\times 10^{-4}$ $5.11\times 10^{-2}$ $7$ $7$ $19$
-$11$ auto $3.5\times 10^{-6}$ 0.0341 0.0347 $2797$ $1.1\times 10^{6}$ $7.91\times 10^{-4}$ $5.04\times 10^{-2}$ $8$ $7$ $20$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+$5$ 0.9 $6.0\times 10^{-5}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $2.95\times 10^{-6}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
+$6$ 0.75 $5.0\times 10^{-5}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $4.30\times 10^{-6}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
+$7$ 0.69 $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.86\times 10^{-6}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
+$8$ 0.66 $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $5.21\times 10^{-6}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
+$9$ 0.61 $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.62\times 10^{-6}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
+$10$ auto $5.0\times 10^{-6}$ 0.0309 0.0357 $198$ $1.0\times 10^{6}$ $6.48\times 10^{-6}$ $5.11\times 10^{-2}$ $7$ $7$ $6$
+$11$ auto $3.5\times 10^{-6}$ 0.0341 0.0347 $270$ $1.1\times 10^{6}$ $8.41\times 10^{-6}$ $5.04\times 10^{-2}$ $8$ $7$ $6$
 (b) $\vec{S}_{0}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{O}$
-$5$ 0.9 $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.37\times 10^{-5}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
-$6$ 0.75 $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.26\times 10^{-5}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
-$7$ 0.64 $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $3.61\times 10^{-5}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
-$8$ 0.57 $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.25\times 10^{-5}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
-$9$ 0.53 $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $5.12\times 10^{-5}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
-$10$ 0.5 $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $6.02\times 10^{-5}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
-$11$ auto $1.0\times 10^{-7}$ 0.0379 0.0401 $853$ $1.1\times 10^{6}$ $3.43\times 10^{-4}$ $5.28\times 10^{-2}$ $9$ $14$ $11$
+$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+$5$ 0.9 $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.24\times 10^{-6}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
+$6$ 0.75 $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.32\times 10^{-6}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
+$7$ 0.64 $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $4.58\times 10^{-6}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
+$8$ 0.57 $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.98\times 10^{-6}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
+$9$ 0.53 $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $7.23\times 10^{-6}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
+$10$ 0.5 $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $7.99\times 10^{-6}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
+$11$ auto $1.0\times 10^{-7}$ 0.0379 0.0401 $125$ $1.1\times 10^{6}$ $6.27\times 10^{-6}$ $5.28\times 10^{-2}$ $9$ $14$ $4$
 Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
@@
 We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs.
-In TT-based option pricing with $\vec{\sigma}$ and $\vec{S_{0}}$, the sharp increase in computational complexity at $d=10$ and $d=11$ arises, respectively.
-This stems from the necessity to set $\eta$ by Eq.
-(19) or (20), which in turn makes the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{m}$ for $\vec{p}$.
 Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$.
@@
 The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
-5.2.2 The case of varying $\vec{S_{0}}$
+V.2.2 The case of varying $\vec{S_{0}}$
 TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence.
@@
 3(b)).
-The maximum bond dimension, $\chi_{O}$, of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is about 6 when $\eta$ is constant, while it increases to 11 when $\eta$ is adjusted per asset, leading to higher computational complexity at about $d=11$.
-Nonetheless, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
+The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 6 or less.
+Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
 Compared to varying $\sigma$, its accuracy is worse at $d=10$.
-5.2.3 Randomness in learning the TTs
+V.2.3 Randomness in learning the TTs
 Here, we mention the randomness in learning the TTs and the error induced by it.
@@
 The aforementioned mean and standard deviation are computed from successful cases.
-5.2.4 Total computational time for obtaining the MPOs
-Lastly, let us mention the total computational time for obtaining the MPOs for $\phi$ and $\hat{v}$ and $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
+V.2.4 Total computational time for obtaining the TTs
+Lastly, let us mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
 We focus on TCI because it dominates over SVD.
@@
 This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec.
-6.
-We also note that, when we do not incorporate the parameter dependence into the tensor networks for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence.
+VI.
+We also note that, when we do not incorporate the parameter dependence into the tensor trains for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence.
 In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1.
 This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
-6 Summary and discussion
+VI Summary and discussion
 We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs.
@@
 This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times.
-However, in our demonstration, compared to the huge computational complexity advantage, the computational time advantage of TT-based option pricing is milder.
-This is because the TT-based method is bottlenecked by specifying the values of $\vec{p}$.
-Specifically, accessing the memory for the specific tensor components $O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ dominates the computational time for additions and multiplications in tensor contraction.
-Therefore, in the TT-based method, and perhaps the MC-based method too, the implementation may have room for improvement to reduce computational time.
+Note that in the MC-based method, the implementation may have room for improvement to reduce computational time.
 We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths.
@@
 An expected way to utilize this method is as follows.
-At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the MPO to quickly price the option for the fluctuating parameters.
+At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the TT to quickly price the option for the fluctuating parameters.
 If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial.
@@
 by increasing $N$ from the current value of 50 to, say, 100.
-It is a promising approach to directly learning the MPOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing.
+It could be a promising approach to directly learning the TTOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing.
 At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself.
@@
 S.
-thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor networks to option pricing.
+thanks the Quantum Software Research Hub at Osaka University for the opportunity to participate in the study group on applications of tensor trains to option pricing.
 R.S.
@@
 JPMJPF2014.
+This work was supported by the Center of Innovation for Sustainable Quantum AI (JST Grant Number JPMJPF2221).
 Appendix A How to set the tolerance $\epsilon_{\mathrm{SVD}}$
@@
 The left vertical axis represents $r_{\mathrm{TT}}$, while the right vertical axis shows $c_{\mathrm{TT}}$.
-Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
-Here, we fixed the tolerance of SVD $\epsilon^{O}_{\mathrm{SVD}}$ to $10^{-10}$.
-The tolerance $\epsilon_{\mathrm{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$, while minimizing the computational complexity $c_{\mathrm{TT}}$.
+Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
+Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-10}$.
+The tolerance $\epsilon_{\mathrm{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$.
 From Fig.
-4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ while minimizing the computational complexity.
+4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$.
 By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.

### v6

@@
 As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices.
-We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ paths in terms of computational complexity while keeping comparable accuracy.
+We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ and $10^{6}$ paths in terms of computational complexity while keeping comparable accuracy.
 I Introduction
@@
 Pricing options appropriately is needed for making a profit and managing the risk of loss in option trading.
-According to the theory in mathematical finance  [^1] As typical textbooks in this area, we refer to [1, 2]., the price of an option is given by the expectation of the discounted payoff in the contract with some stochastic model on the dynamics of underlying asset prices assumed.
+According to the theory in mathematical finance  [^1] As typical textbooks in this area, we refer to Refs.
+[1, 2]., the price of an option is given by the expectation of the discounted payoff in the contract with some stochastic model on the dynamics of underlying asset prices assumed.
 Except for limited cases with simple contract conditions and models, the analytical formula for the option price is not available and thus we need to resort to the numerical calculation.
@@
 Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
-Motivated by these points, recently, applications of quantum computing to option pricing are considered actively [^2] See [3] as a comprehensive review..
+Motivated by these points, recently, applications of quantum computing to option pricing are considered actively [^2] See Ref.
+[3] as a comprehensive review..
 For example, many studies have focused on applications of the quantum algorithm for Monte Carlo integration [4], which provides the quadratic quantum speed-up over the classical counterpart.
 Unfortunately, running such a quantum algorithm requires fault-tolerant quantum computers, which may take decades to be developed.
-In light of this, applications of quantum-inspired algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present or the near future [5, 6, 7, 8].
+In light of this, applications of quantum-inspired classical algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present  [5, 6, 7, 8].
 Among them, this paper focuses on the application of tensor train (TT) [9]  [^3] Tensor train is a mathematically equivalent to matrix product state (MPS) [10].
-learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in [6].
+learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in Ref.
+[6].
 This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$.
@@
 Here, we would like to point out an issue in this TT-based method.
-That is, we need to rerun TCI to obtain TTs each time the parameters are changed.
+That is, we need to rerun TCI to obtain TTs each time the input parameters of option price are changed.
 According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
-To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [28, 29, 30] to make FT-based option pricing more efficient.
+To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [28, 29, 30, 31] to make FT-based option pricing more efficient.
 Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function.
-We use these tensor trains to perform fast option pricing in response to various parameter changes.
+We use these tensor trains to evaluate the integral including parameter dependence and perform fast option pricing in response to various parameter changes (refer to FIG.
+1).
 To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices.
-In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice.
-Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ up to $d=11$ (refer to FIG.
+In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ and $10^{6}$ paths by a factor of $O(10^{4})$ and $O(10^{5})$, respectively (see Fig.
 2).
-We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths.
+We also confirm numerically that in the tested cases, the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths, and, although worse in some cases, is overall comparable to that with $10^{6}$ paths.
 In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
+In the context of TT-based approximations of high-dimensional functions, incorporating parameter dependence into TTs has been considered in some fields [28, 29, 30, 31].
+However, to the best of our knowledge, our study is the first to take such an approach in FT-based option pricing in order to make it more efficient for varying parameters, which provides practical benefits in the rapidly changing financial market.
 The rest of this paper is organized as follows.
@@
 In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low.
-This is done by first canonicalizing the TT using QR decompositions, and then performing the compression via SVD, discarding singular values that are smaller than the threshold $\epsilon_{\mathrm{SVD}}$ set for each bond.
+This is done by first canonicalizing the TT using QR decompositions, and then performing the compression via SVD, discarding singular values that are smaller than the tolerance $\epsilon_{\mathrm{SVD}}$ set for each bond.
+This tolerance is defined by
+$\epsilon_{\mathrm{SVD}}=\frac{|\tilde{F}_{\mathrm{TT}}-\tilde{F}^{{}^{\prime}}_{\mathrm{TT}}|^{2}_{\mathrm{F}}}{|\tilde{F}_{\mathrm{TT}}|^{2}_{\mathrm{F}}},$ (4)
+where $|\cdots|^{2}_{\mathrm{F}}$ indicates the Frobenius norm, $\tilde{F}_{\mathrm{TT}}$ is the TT obtained from TCI and $\tilde{F}^{{}^{\prime}}_{\mathrm{TT}}$ is the other TT after SVD.
 For more technical details, readers are referred to Ref.
@@
 samples from $q(\vec{x}|\vec{x}_{0})$.
-On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., [32]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions.
+On how to sample multivariate normal variables, we leave the detail to textbooks (e.g., Ref.
+[33]) and just mention that it requires more complicated operations than simple multiplications and additions such as evaluations of some elementary functions.
 Besides, calculating the payoff $v$ with the normal variable $\vec{x}_{i}$ involves exponentiation.
@@
 We reduce the bond dimensions of these TTs using SVD.
-We then contract adjacent core tensors pairwise and obtain two tensor train operators (TTOs), also known as matrix product operator (MPO), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$.
+We then contract adjacent core tensors pairwise and obtain two tensor train operators (TTOs), also known as matrix product operators (MPOs), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$.
 We contract these TTOs along the index vector $\vec{j}$ as follows:
@@
 IV.2 Computational complexity
-[…]
+The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $\tilde{V}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{\tilde{V}}^{2})$ [10].
+Here, we denote the maximum bond dimensions as $\chi_{V}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations.
+Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec.
+V.
 Here, we ignore the computational complexity of all the processes in FIG.
@@
 As discussed in Sec.
-VI, we can reasonably find such a situation in practice.
+VII, we can reasonably find such a situation in practice.
 V Numerical demonstration
@@
 We take the MC-based method with $10^{5}$ paths, which is a typical number in option pricing in practical business, as the comparison target against our method.
-The average relative error and its standard deviation of the MC-based method with $10^{5}$ paths are calculated from its 20 runs.
+We also run the MC method with $10^{6}$ paths, as the value when high accuracy is needed.
+As the error of the MC-based option price, we compute the half-width of its 95% confidence interval, namely $1.96\sigma/\sqrt{n_{\rm MC}}$, where $\sigma$ is the sample standard deviation of the discounted payoff in the MC run and the path number $n_{\rm MC}$ is $10^{5}$ or $10^{6}$.
 We assess the accuracy of our method by seeing if it falls within the statistical error range of the MC-based method.
@@
 Thus, we randomly select 100 combinations and perform option pricing for each of them.
-We compare the maximum among the relative errors of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
+We compare the maximum among the errors of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
 Software and hardware used in this study
-TensorCrossInterpolation.jl for TCI [34] has not yet been made publicly available.
-Marc K.
-Ritter, Hiroshi Shinaoka, and Jan von Delft provided us with access to this.
+TensorCrossInterpolation.jl [35] was used for learning tensor trains from functions appearing Fourier-based option pricing.
 ITensors.jl [36] was utilized for computing the contraction and inner product of the tensor trains.
@@
 The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
-V.2 Results
+VI Results
 We show the results for the computational complexity, time, and accuracy of TT-based and MC-based option pricing when two parameters $\vec{\sigma}$ and $\vec{S_{0}}$ are varied.
@@
 2.
-The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$.
-The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively.
-The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively.
-Also, the computational time of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively.
+The maximum error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$.
+For the parameter set that yields this maximum error, we compute the half-width of the 95% confidence interval for the MC-based option price and denote it by $e_{\mathrm{MC},10^{5}}$ and $e_{\mathrm{MC},10^{6}}$ for the cases with $10^{5}$ and $10^{6}$ paths, respectively.
+The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{5}$ paths are denoted by $c_{\mathrm{MC},10^{5}}$ and $t_{\mathrm{MC},10^{5}}$, with those for $10^{6}$ paths being 10 times larger.
 To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$.
-Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
+Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
 The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2.
-We set the tolerance of SVD for obtaining $\chi_{V}$, maximum bond dimension of $\tilde{V}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
-V.2.1 The case of varying $\vec{\sigma}$
+In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
+VI.0.1 The case of varying $\vec{\sigma}$
 TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence.
@@
 3(a).
-By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at less than 10.
-The details of this compression by SVD are described in Appendix A.
-The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of 6 or less, regardless of when $\eta$ is constant or $\eta$ is automatically determined for each asset.
+By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10 when $\eta$ is constant.
+When $\eta$ is automatically determined for each asset, the bond dimensions are around 60 or 70.
+The details of this compression by SVD are described in Appendix.
+The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, regardless of when $\eta$ is constant or $\eta$ is automatically determined for each asset.
 Therefore, the computational complexity is much lower than that of MC-based option pricing.
-The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.
-It has been observed that accuracy tends to decrease as $d$ increases.
-Consequently, increasing $d$ further might result in decreased accuracy, making it difficult to maintain accuracy within the MC-based method.
+We compare the TT-based results with the MC simulations with $10^{5}$ and $10^{6}$ paths.
+With $10^{5}$ paths, the TT-based option price is within the MC error range in all the tested cases.
+With $10^{6}$ paths, the TT-based method overall gives the errors comparable with the MC-based method, though it performes worse for $d=8$ and $d=9$.
 Furthermore, reducing the number of subdivisions below 100 resulted in only minor changes in accuracy and computational complexity.
 (a) $\vec{\sigma}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
-$5$ 0.9 $6.0\times 10^{-5}$ 0.0162 0.0207 $43$ $5.0\times 10^{5}$ $2.95\times 10^{-6}$ $2.84\times 10^{-2}$ $4$ $7$ $4$
-$6$ 0.75 $5.0\times 10^{-5}$ 0.0231 0.0254 $75$ $6.0\times 10^{5}$ $4.30\times 10^{-6}$ $3.38\times 10^{-2}$ $5$ $7$ $5$
-$7$ 0.69 $1.8\times 10^{-5}$ 0.0203 0.0236 $70$ $7.0\times 10^{5}$ $4.86\times 10^{-6}$ $3.71\times 10^{-2}$ $6$ $7$ $4$
-$8$ 0.66 $2.0\times 10^{-5}$ 0.0179 0.0253 $73$ $8.0\times 10^{5}$ $5.21\times 10^{-6}$ $4.09\times 10^{-2}$ $6$ $7$ $4$
-$9$ 0.61 $1.0\times 10^{-5}$ 0.0240 0.0258 $89$ $9.0\times 10^{5}$ $5.62\times 10^{-6}$ $4.51\times 10^{-2}$ $7$ $7$ $4$
-$10$ auto $5.0\times 10^{-6}$ 0.0309 0.0357 $198$ $1.0\times 10^{6}$ $6.48\times 10^{-6}$ $5.11\times 10^{-2}$ $7$ $7$ $6$
-$11$ auto $3.5\times 10^{-6}$ 0.0341 0.0347 $270$ $1.1\times 10^{6}$ $8.41\times 10^{-6}$ $5.04\times 10^{-2}$ $8$ $7$ $6$
+$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+$5$ 0.9 $1.0\times 10^{-5}$ 0.00560 0.0213 0.00670 $14$ $5.0\times 10^{5}$ $7.77\times 10^{-7}$ $2.84\times 10^{-2}$ 12 9 2
+$6$ 0.75 $1.0\times 10^{-5}$ 0.00377 0.0143 0.00456 $18$ $6.0\times 10^{5}$ $9.37\times 10^{-7}$ $3.38\times 10^{-2}$ 13 9 2
+$7$ 0.69 $1.0\times 10^{-5}$ 0.00380 0.0124 0.00382 $22$ $7.0\times 10^{5}$ $9.12\times 10^{-7}$ $3.71\times 10^{-2}$ 13 8 2
+$8$ 0.66 $1.5\times 10^{-5}$ 0.00547 0.0128 0.00346 $26$ $8.0\times 10^{5}$ $1.75\times 10^{-6}$ $4.09\times 10^{-2}$ 14 8 2
+$9$ 0.61 $1.5\times 10^{-5}$ 0.00694 0.0100 0.00302 $30$ $9.0\times 10^{5}$ $2.47\times 10^{-6}$ $4.51\times 10^{-2}$ 13 7 2
+$10$ auto $1.5\times 10^{-6}$ 0.00230 0.00916 0.00282 $34$ $1.0\times 10^{6}$ $2.00\times 10^{-6}$ $5.11\times 10^{-2}$ 59 58 2
+$11$ auto $1.5\times 10^{-7}$ 0.00117 0.00847 0.00269 $38$ $1.1\times 10^{7}$ $2.16\times 10^{-6}$ $5.04\times 10^{-2}$ 73 76 2
 (b) $\vec{S}_{0}$
-$d$ $\eta$ $\epsilon_{\mathrm{SVD}}$ $r_{\mathrm{TT}}$ $r_{\mathrm{MC}}+\sigma_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[sec] $t_{\mathrm{MC}}$[sec] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
-$5$ 0.9 $2.0\times 10^{-4}$ 0.0328 0.0374 $81$ $5.0\times 10^{5}$ $3.24\times 10^{-6}$ $2.90\times 10^{-2}$ $4$ $6$ $6$
-$6$ 0.75 $7.0\times 10^{-5}$ 0.0265 0.0279 $114$ $6.0\times 10^{5}$ $4.32\times 10^{-6}$ $3.28\times 10^{-2}$ $5$ $6$ $6$
-$7$ 0.64 $3.0\times 10^{-5}$ 0.0308 0.0355 $139$ $7.0\times 10^{5}$ $4.58\times 10^{-6}$ $3.66\times 10^{-2}$ $6$ $7$ $6$
-$8$ 0.57 $1.0\times 10^{-5}$ 0.0259 0.0340 $120$ $8.0\times 10^{5}$ $5.98\times 10^{-6}$ $3.93\times 10^{-2}$ $7$ $7$ $5$
-$9$ 0.53 $1.0\times 10^{-5}$ 0.0311 0.0345 $163$ $9.0\times 10^{5}$ $7.23\times 10^{-6}$ $4.54\times 10^{-2}$ $7$ $7$ $5$
-$10$ 0.5 $3.0\times 10^{-6}$ 0.0315 0.0414 $171$ $1.0\times 10^{6}$ $7.99\times 10^{-6}$ $4.93\times 10^{-2}$ $7$ $8$ $5$
-$11$ auto $1.0\times 10^{-7}$ 0.0379 0.0401 $125$ $1.1\times 10^{6}$ $6.27\times 10^{-6}$ $5.28\times 10^{-2}$ $9$ $14$ $4$
+$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+$5$ 0.9 $1.0\times 10^{-5}$ 0.0195 0.0420 0.0133 $14$ $5.0\times 10^{5}$ $7.67\times 10^{-7}$ $3.99\times 10^{-2}$ 14 9 2
+$6$ 0.75 $1.0\times 10^{-5}$ 0.0165 0.0385 0.0121 $18$ $6.0\times 10^{5}$ $9.34\times 10^{-7}$ $5.15\times 10^{-2}$ 16 9 2
+$7$ 0.69 $3.0\times 10^{-5}$ 0.0103 0.0357 0.0112 $22$ $7.0\times 10^{5}$ $9.21\times 10^{-7}$ $5.37\times 10^{-2}$ $21$ $10$ $2$
+$8$ 0.57 $3.5\times 10^{-5}$ 0.0125 0.0303 0.00952 $26$ $8.0\times 10^{5}$ $1.63\times 10^{-6}$ $6.18\times 10^{-2}$ $17$ $7$ $2$
+$9$ 0.53 $1.0\times 10^{-5}$ 0.0155 0.0204 0.00650 $30$ $9.0\times 10^{5}$ $1.19\times 10^{-6}$ $7.95\times 10^{-2}$ $19$ $7$ $2$
+$10$ 0.5 $3.0\times 10^{-5}$ 0.0172 0.0204 0.00654 $34$ $1.0\times 10^{6}$ $1.47\times 10^{-6}$ $8.77\times 10^{-2}$ $16$ $6$ $2$
+$11$ auto $1.0\times 10^{-7}$ 0.0139 0.0181 0.00579 $38$ $1.1\times 10^{6}$ $1.93\times 10^{-6}$ $8.43\times 10^{-2}$ $48$ $38$ $2$
 Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
 Here, we set the ranges $\sigma_{m}\in[0.15,0.25]$ and $S_{m,0}\in[90,120]$, and place 100 equally spaced grid points within these ranges.
-The $\epsilon_{\mathrm{SVD}}$ in the table refers to selected values that reduce $c_{\mathrm{TT}}$ keeping $r_{\mathrm{TT}}$ within the error range of MC-based option pricing $r_{\mathrm{MC}}+\sigma_{\text{MC}}$.
+The $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in the table refers to the values adjusted so that $c_{\mathrm{TT}}$ is reduced with $e_{\mathrm{TT}}$ kept smaller than the MC error $e_{\text{MC},10^{5}}$.
 The term auto in (a) and (b) indicate that $\eta_{m}$ was automatically determined for each asset based on Eqs.
@@
 $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
-Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, versus the number of assets $d$.
-We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs.
+Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$ versus the number of assets $d$.
+We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{5}$ and $10^{6}$.
 Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$.
@@
 The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
-V.2.2 The case of varying $\vec{S_{0}}$
+VI.0.2 The case of varying $\vec{S_{0}}$
 TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence.
@@
 3(b)).
-The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 6 or less.
-Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
-Compared to varying $\sigma$, its accuracy is worse at $d=10$.
-V.2.3 Randomness in learning the TTs
+For $d=11$, when $\eta$ is determined automatically, this bond dimension increases to around 40.
+The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 2 Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
+Compared with the MC-based method with $10^{5}$ paths, the TT-based mathod again gives the smaller error in every value of $d$.
+With $10^{6}$ paths, the TT error is worse than the MC error for many values of $d$, but still comparable.
+VI.0.3 Randomness in learning the TTs
 Here, we mention the randomness in learning the TTs and the error induced by it.
@@
 The aforementioned mean and standard deviation are computed from successful cases.
-V.2.4 Total computational time for obtaining the TTs
-Lastly, let us mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
+VI.0.4 Total computational time for obtaining the TTs
+We mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
 We focus on TCI because it dominates over SVD.
@@
 This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec.
-VI.
-We also note that, when we do not incorporate the parameter dependence into the tensor trains for $\phi$ and $\hat{v}$ as in [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence.
+VII.
+We also note that, when we do not incorporate the parameter dependence into the tensor trains for $\phi$ and $\hat{v}$ as in Ref.
+[6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence.
 In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1.
 This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
-VI Summary and discussion
+VII Discussion
 We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs.
@@
 Up to $d=11$, we demonstrated superiority in both computational complexity and time.
-Here, we argue that, in principle, given at least comparable computational complexity, TT-based option pricing has an advantage in terms of computational time.
-This is because the TT-based method involves only multiplication and addition as unit operations, while the MC-based method requires more complex operations, such as random number generation and exponentiation, to be iterated $dN_{\rm path}$ times.
 Note that in the MC-based method, the implementation may have room for improvement to reduce computational time.
-We also numerically verified that all TT-based results fell within the statistical error range of the Monte Carlo method with $10^{5}$ paths.
-However, it should be noted that increasing $d$ worsens accuracy, indicating a tendency for the accuracy to be worse than that of the MC-based method around $d\geq 12$.
+We numerically verified that all TT-based results fell within the statistical error range of the MC-based method with $10^{5}$ paths.
+Compared to the MC-based method with $10^{6}$ paths, the TT-based method gave the worse errors in some cases, but was still comparable.
+To enhance the accuracy of the TT-based method, we may reduce the tolerance of the TCI further or change the initial randomly selected grid points of TCI.
 Now, let us consider how the proposed method provides benefits in practical business in financial firms.
@@
 For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step.
-Expanding the methodology to other types of option calculations, particularly the application to American options, is also an important research direction.
-Acknowledgements
+Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing.
+With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [38] and stochastic volatility model [39, 40] are also used in practice.
+Applying our method to such models, where the dependencies on the parameters in the models are incorporated into TTs, is an interesting direction.
+Expanding the scope to the broad types of products is also desired.
+While this paper focused on European-type options, more complicated ones such as American and Bermudan options are also traded widely, and pricing of such options is more time-consuming.
+Considering whether our method can be extended for such options will be an important and interesting challenge.
+VIII Acknowledgements
 We are grateful to Marc K.
@@
 This work was supported by the Center of Innovation for Sustainable Quantum AI (JST Grant Number JPMJPF2221).
-Appendix A How to set the tolerance $\epsilon_{\mathrm{SVD}}$
-Figure 4: The maximum relative error $r_{\mathrm{TT}}$ and computational complexity $c_{\mathrm{TT}}$ when applying SVD with various $\epsilon_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
-The left vertical axis represents $r_{\mathrm{TT}}$, while the right vertical axis shows $c_{\mathrm{TT}}$.
-Figure 4 shows the maximum relative error $r_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
-Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-10}$.
-The tolerance $\epsilon_{\mathrm{SVD}}$ for SVD was chosen to maintain $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$.
+IX Author contributions statement
+R.S.
+conceived the project, and R.S.
+and H.T.
+wrote the code with help from K.M.
+and performed numerical simulations.
+All authors actively discussed the results and contributed to the preparation of the manuscript.
+X Appendix
+X.1 How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
+Figure 4: The maximum error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
+Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
+Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-9}$.
+The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$.
 From FIG.
-4, it can be seen that $r_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-5}$ is appropriate for keeping $r_{\mathrm{TT}}$ smaller than $r_{\mathrm{MC}}+\sigma_{\text{MC}}$.
+4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$.
 By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
-From the fact that we can keep the maximum relative error $r_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information.
+From the fact that we can keep the maximum relative error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information.
 By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed.

### v7

@@
 As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices.
-We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ and $10^{6}$ paths in terms of computational complexity while keeping comparable accuracy.
+We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{6}$ paths in terms of computational complexity while keeping better accuracy.
 I Introduction
@@
 Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
+The quasi-Monte Carlo method has a better convergence rate in low-dimensional cases, but it can deteriorate in high-dimensional cases.
 Motivated by these points, recently, applications of quantum computing to option pricing are considered actively [^2] See Ref.
@@
 Ref.
-[6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing.
+[6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing  [^5] As another approach for accelerating FT-based multi-asset option pricing, see [25]..
 The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [26, 27, 19, 28] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
 To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices.
-In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ and $10^{6}$ paths by a factor of $O(10^{4})$ and $O(10^{5})$, respectively (see Fig.
+In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{6}$ paths by a factor of $O(10^{5})$ (see Fig.
 2).
-We also confirm numerically that in the tested cases, the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths, and, although worse in some cases, is overall comparable to that with $10^{6}$ paths.
-In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
+We also confirm numerically that in the tested cases, the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{6}$ paths.
+In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the better accuracy.
 In the context of TT-based approximations of high-dimensional functions, incorporating parameter dependence into TTs has been considered in some fields [29, 30, 31, 32].
@@
 Tensor cross interpolation (TCI) is a technique to compress tensors corresponding to discretized multivariate functions with a low-rank TT representation.
-Here, we consider a tensor that, with grid points set in $\mathbb{R}^{d}$, has entries $F_{x_{1},\ldots,x_{d}}$ equal to $F(x_{1},\ldots,x_{d})$, the values of a function $F$ on the grid points [^5] Although we here denote the indexed of the tensor and the variables of the function by the same symbols $x_{1},\ldots,x_{d}$ for illustrative presentation, we assume that, in reality, the grid points in $\mathbb{R}^{d}$ is labeled by integers and the indexes of the tensor denotes the integers..
+Here, we consider a tensor that, with grid points set in $\mathbb{R}^{d}$, has entries $F_{x_{1},\ldots,x_{d}}$ equal to $F(x_{1},\ldots,x_{d})$, the values of a function $F$ on the grid points [^6] Although we here denote the indexes of the tensor and the variables of the function by the same symbols $x_{1},\ldots,x_{d}$ for illustrative presentation, we assume that, in reality, the grid points in $\mathbb{R}^{d}$ is labeled by integers and the indexes of the tensor denotes the integers..
 Leaving the detailed explanation to Refs.
@@
 $V(\vec{p})=\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\phi(-\vec{z}_{\mathrm{gr},\vec{j}}-i\vec{\alpha})\hat{v}(\vec{z}_{\mathrm{gr},\vec{j}}+i\vec{\alpha})\Delta_{\rm vol}.$ (13)
-Here, the even natural number $N$ is the number of the grid points in one dimension. $\vec{z}_{\mathrm{gr},\vec{j}}$ is the grid point specified by the integer vector $\vec{j}$ as
-$\vec{z}_{\mathrm{gr},\vec{j}}:=(\eta_{1}(\vec{p})j_{1},\ldots,\eta_{d}(\vec{p})j_{d}),$ (14)
-where $\eta_{m}$ is the integration step size in the $m$-th direction.
-$\eta_{m}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{m}(\vec{p})$ in (14).
-In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec.
-V for the detail.
-$\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
+Here, the even natural number $N$ is the number of the grid points in one dimension,$\vec{z}_{\mathrm{gr},\vec{j}}$ is the grid point specified by the integer vector $\vec{j}$ as
+$\vec{z}_{\mathrm{gr},\vec{j}}:=(\eta j_{1},\ldots,\eta j_{d}),$ (14)
+where $\eta$ is the integration step size in each direction, and $\Delta_{\rm vol}:=\eta^{d}$ is the volume element.
+$\eta$ is a hyperparameter that must be appropriately determined.
+In our demonstration, it is set to 0.4 or 0.3, which yields the accurate result (see Sec.
+V for details).
 III.2 Fourier transform-based option pricing with tensor trains
@@
 Then, extending the aforementioned FT-based option pricing method with TTs, we propose a new scheme to quickly compute the option price in response to the change in the input parameter set.
-Using TCI, we obtain the TTs to approximate $\phi$ and $\hat{v}$, incorporating the parameter dependence of these functions.
+Using TCI, we obtain the TTs to approximate $\phi$ incorporating the parameter dependence of this function and $\hat{v}$.
 The outline is illustrated in FIG.
@@
 Here, focusing not all the parameters but a part of them, we take $\vec{p}$ as a $d$-dimensional vector, e.g., either of $\vec{\sigma}$ or $\vec{S}_{0}$.
-Considering $\phi$ and $\hat{v}$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$.
+Considering $\phi$ as the functions of not only $\vec{z}$ but also $\vec{p}$, we set in the space of $\vec{z}$ and $\vec{p}$ the grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ labeled by the index vectors $\vec{j}$ and $\vec{k}$.
 Then, as illustrated in FIG.
-1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$.
+1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},\ldots,j_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ and $(\vec{z}_{\mathrm{gr},\vec{j}})$, respectively.
 We reduce the bond dimensions of these TTs using SVD.
-We then contract adjacent core tensors pairwise and obtain two tensor train operators (TTOs), also known as matrix product operators (MPOs), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and $\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}}$.
-We contract these TTOs along the index vector $\vec{j}$ as follows:
-$\tilde{V}^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
-Note that this resultant TTO has only the diagonal elements with respect to the local indices $\vec{k}$.
-That is, when $k_{1}\neq k^{\prime}_{1},\ldots,k_{d}\neq k^{\prime}_{d}$, we set $\tilde{V}^{k_{1},\ldots,k_{d}}_{k^{\prime}_{1},\ldots,k^{\prime}_{d}}=0$.
-With this operation, the resultant tensor can be expressed as TT $\tilde{V}_{k_{1},\ldots,k_{d}}$.
+We then contract adjacent core tensors pairwise and obtain a tensor train operators (TTO), also known as matrix product operators (MPO), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$.
+We contract this TTO $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$.
+and TT $\tilde{v}_{j_{1},\ldots,j_{d}}$ along the index vector $\vec{j}$ as follows:
+$\tilde{V}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}},$ (17)
 Then we optimize the bond dimensions of $\tilde{V}_{k_{1},\ldots,k_{d}}$ using SVD.
+Thus, we get a new function $\tilde{V}_{k_{1},\ldots,k_{d}}$ whose input are parameter and output is option price.
 Having this TT, we get the option price for the specified value of $\vec{p}$ as illustrated in FIG.
@@
 $V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}\tilde{V}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
-Here, we note that the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ for $\hat{v}$ and its TT approximation $\tilde{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ have indexes $k_{i}$ corresponding to $\vec{p}$, which is now $\vec{\sigma}$ or $\vec{S}_{0}$, the parameter not affecting $\hat{v}$.
-This is due to $\eta_{j}$ depending on $\vec{p}$: because of the change of the grid points in the $\vec{z}$ space with respect to $\vec{p}$, the tensor $\hat{v}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ depends on the indexes $k_{i}$.
-We also mention the ordering of the local indices of the TT.
+Here, we mention the ordering of the local indices of the TT.
 Two core tensors in the TT that correspond to $z_{j}$ and $p_{j}$ of the same asset are arranged next to each other, i.e., the order is ($z_{1}p_{1}z_{2}p_{2}\cdots z_{d}p_{d}$).
@@
 Figure 1: Fast option pricing based on TNs proposed in this study.
-In (a), we learn TTs with the parameter dependence of the functions using TCI and reduce the bond dimension of these TTs using SVD.
-We then contract the tensors associated with $p_{m}$ and $z_{m}$, resulting in two TTOs.
-We then take the contraction of these two TTOs with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting TT using SVD.
+In (a), we learn TTs from parameter-dependent function of $\phi$ and $\hat{v}$ using TCI and reduce the bond dimension of these TTs using SVD.
+We then contract the tensors associated with $p_{m}$ and $z_{m}$ for TT of $\phi$, resulting in TTO.
+We then take the contraction of the TTO and TT with respect to the index vector $\vec{j}$ and compress the bond dimension of the resulting TT using SVD.
 In (b), we use this TT to perform fast option pricing for a specified parameter $\vec{p}$.
@@
 In the proposed method, we need to set the range in the space of $\vec{p}$ and the grid points in it.
-For each volatility $\sigma_{m}$, we set the range to $\sigma_{m}\in[0.15,0.25]$, where the center $\sigma_{m}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [34] and the range width ±$\pm 0.05$ covers the changes in this index on most days.
-For each $S_{m,0}$, we set the range to $S_{m,0}\in[90,120]$, which corresponds to the 20% variation of the asset price centered at 100.
+For each volatility $\sigma_{m}$, we set the range to $\sigma_{m}\in[0.15,0.25)$, where the center $\sigma_{m}=0.2$ is a typical value of Nikkei Stock Average Volatility Index [35] and the range width ±$\pm 0.05$ covers the changes in this index on most days.
+For each $S_{m,0}$, we set the range to $S_{m,0}\in[90,120)$, which corresponds to the 20% variation of the asset price centered at 100.
 The lower bound is set to not 80 but 90 because the price of the option we take as an example is negligibly small for $S_{0}<90$.
@@
 The other parameters for option pricing are fixed to the values summarized in the table below.
-$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{mm}$ $\rho_{mn(m\neq n)}$ $N$
-$1$ $0.01$ $100$ $100$ $5/d$ $1$ $1/3$ $50$
+$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{mn},m\neq n$ $(N,\eta)$
+$1$ $0.01$ $100$ $100$ $\frac{5}{d}$ $\frac{1}{3}$$\left\{\begin{array}[]{ll}(200,0.3)&;\ d=11,\vec{p}=\vec{S}_{0}\\ (100,0.4)&;\ \mathrm{otherwise}\end{array}\right.$
 Table 1: The parameters except $\vec{\sigma}$ and $\vec{S}_{0}$.
-Integral step size
-The spread of the characteristic function changes depending on the values of the parameters $\vec{p}$ and the number of assets.
-Accordingly, it is necessary to adjust the grid interval $\eta_{m}$ for integration in the wavenumber space.
-For example, consider the case where $\vec{p}=\vec{\sigma}$.
-From Eq.
-(10), we see that for the current parameter values in TABLE 1, $\phi$ takes non-negligible value only when $\sigma_{m}z_{m}\lesssim 1$ for each $m$.
-It is necessary to ensure that the grid points cover the whole of such a region with a sufficiently fine interval.
-Currently, as $z_{m}$ ranges from $-N\eta$ to $+N\eta$ with step size $\eta_{m}$, $\eta_{m}$ and $\sigma_{m}$ must satisfy $\sigma_{m}\eta_{m}\ll 1$ and $\sigma_{m}\times N\eta_{m}\gg 1$.
-Therefore, in general, we need the adjustment that $\eta_{m}$ increases when $\sigma$ decreases and vice versa.
-Nevertheless, in almost all the cases we tested, we observed that we get the accurate option price with $\eta_{m}$ set to some constant, and thus we did so.
-In the cases with $d=10,11$, $\eta_{m}$ is adjusted as
-$\eta_{m}=\frac{a}{\sigma_{m}}+b$ (19)
-To set the constants $a$ and $b$, we ran TT-based option pricing for several values of $\sigma_{m}$ adjusting $\eta_{m}$, found the pairs $(\sigma_{m},\eta_{m})$ that yields the accurate price, and adjust $a$ and $b$ so that Eq.
-(19) approximately reproduces such $\eta_{m}$ from $\sigma_{m}$.
-Also when $\vec{p}=\vec{S}_{0}$, we set $\eta_{m}$ to some constant in almost all the test cases. Only for $d=11$, we try the adjustment that
-$\eta_{m}=\frac{\alpha}{S_{m,0}}+\beta$ (20)
-with $\alpha$ and $\beta$ chosen according to the option pricing result for several $\eta_{m}$ and $\sigma_{m}$.
-In general, larger $S_{m,0}$, which means larger $x_{m,0}$, makes $\phi$ as a function of $z_{m}$ oscillate more rapidly, and thus we need finer grid points to resolve such an oscillation.
-Note that when $\eta_{m}$’s are set constant, the grid points do not depend on $\vec{p}$, and thus we can take the tensor $\hat{v}_{j_{1},\ldots,j_{d}}$ and its TT approximation $\tilde{v}_{j_{1},\ldots,j_{d}}$ having only the indexes $j_{1},\ldots,j_{d}$ associated with $z_{1},\ldots,z_{d}$.
 Error evaluation
@@
 Instead, we regard the option price computed by the MC-based method with very many paths, concretely $5\times 10^{7}$ paths, as the true value.
-The error of the option price computed by the proposed method is evaluated by its deviation from the true value.
-We take the MC-based method with $10^{5}$ paths, which is a typical number in option pricing in practical business, as the comparison target against our method.
-We also run the MC method with $10^{6}$ paths, as the value when high accuracy is needed.
-As the error of the MC-based option price, we compute the half-width of its 95% confidence interval, namely $1.96\sigma/\sqrt{n_{\rm MC}}$, where $\sigma$ is the sample standard deviation of the discounted payoff in the MC run and the path number $n_{\rm MC}$ is $10^{5}$ or $10^{6}$.
-We assess the accuracy of our method by seeing if it falls within the statistical error range of the MC-based method.
-An issue is that the number of possible combinations of the parameter $\vec{p}$ is $100^{d}$, and thus we cannot test all of them.
+The error of the option price computed by the proposed method is evaluated by calculating the mean absolute error from the true value.
+As a comparison target for our method, we use an MC-based approach with $10^{6}$ paths.
+As the error of the MC-based option price, we compute the half-width of its 95% confidence interval, namely $1.96\sigma/\sqrt{n_{\rm MC}}$, where $\sigma$ is the sample standard deviation of the discounted payoff in the MC run and the path number $n_{\rm MC}$ is $10^{6}$.
+We assess the accuracy of our method by verifying if its error is below that of the MC-based method.
+Here, an issue is that the number of possible combinations of the parameter $\vec{p}$ is $100^{d}$, and thus we cannot test all of them.
 Thus, we randomly select 100 combinations and perform option pricing for each of them.
-We compare the maximum among the errors of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
+We compare the mean absolute error of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
 Software and hardware used in this study
@@
 GPUs were not utilized in any of the calculations.
-The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
+The computations were performed on a 2023 MacBook Pro featuring a 12-core Apple M2 Max processor and 32GB of 400GB/s unified memory.
 VI Results
@@
 2.
-The maximum error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$.
-For the parameter set that yields this maximum error, we compute the half-width of the 95% confidence interval for the MC-based option price and denote it by $e_{\mathrm{MC},10^{5}}$ and $e_{\mathrm{MC},10^{6}}$ for the cases with $10^{5}$ and $10^{6}$ paths, respectively.
-The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{5}$ paths are denoted by $c_{\mathrm{MC},10^{5}}$ and $t_{\mathrm{MC},10^{5}}$, with those for $10^{6}$ paths being 10 times larger.
-To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$.
-Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively.
+The mean absolute error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$.
+For the same parameter sets, we compute mean values of the absolute values of the error in the MC-based option price with $10^{6}$ paths and denote it by $e_{\mathrm{MC}}$.
+The computational complexity and time of TT-based option pricing are represented by $c_{\mathrm{TT}}$ and $t_{\mathrm{TT}}$, respectively, and those of MC-based pricing with $10^{6}$ paths are denoted by $c_{\mathrm{MC}}$ and $t_{\mathrm{MC}}$.
+To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$.
+Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}=1.0\times 10^{-6}$ for $\phi$ and $\hat{v}$.
 The maximum bond dimensions of the TTs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2.
-In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
+In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=1.0\times 10^{-6}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
 VI.0.1 The case of varying $\vec{\sigma}$
@@
 By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10 when $\eta$ is constant.
-When $\eta$ is automatically determined for each asset, the bond dimensions are around 60 or 70.
 The details of this compression by SVD are described in Appendix.
-The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, regardless of when $\eta$ is constant or $\eta$ is automatically determined for each asset.
+A.
+The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, or 1 especially when $d=10,11$.
 Therefore, the computational complexity is much lower than that of MC-based option pricing.
-We compare the TT-based results with the MC simulations with $10^{5}$ and $10^{6}$ paths.
-With $10^{5}$ paths, the TT-based option price is within the MC error range in all the tested cases.
-With $10^{6}$ paths, the TT-based method overall gives the errors comparable with the MC-based method, though it performes worse for $d=8$ and $d=9$.
-Furthermore, reducing the number of subdivisions below 100 resulted in only minor changes in accuracy and computational complexity.
+Besides, the error in the TT-based result is smaller than that in the MC-based result in all the tested cases.
 (a) $\vec{\sigma}$
-$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
-$5$ 0.9 $1.0\times 10^{-5}$ 0.00560 0.0213 0.00670 $14$ $5.0\times 10^{5}$ $7.77\times 10^{-7}$ $2.84\times 10^{-2}$ 12 9 2
-$6$ 0.75 $1.0\times 10^{-5}$ 0.00377 0.0143 0.00456 $18$ $6.0\times 10^{5}$ $9.37\times 10^{-7}$ $3.38\times 10^{-2}$ 13 9 2
-$7$ 0.69 $1.0\times 10^{-5}$ 0.00380 0.0124 0.00382 $22$ $7.0\times 10^{5}$ $9.12\times 10^{-7}$ $3.71\times 10^{-2}$ 13 8 2
-$8$ 0.66 $1.5\times 10^{-5}$ 0.00547 0.0128 0.00346 $26$ $8.0\times 10^{5}$ $1.75\times 10^{-6}$ $4.09\times 10^{-2}$ 14 8 2
-$9$ 0.61 $1.5\times 10^{-5}$ 0.00694 0.0100 0.00302 $30$ $9.0\times 10^{5}$ $2.47\times 10^{-6}$ $4.51\times 10^{-2}$ 13 7 2
-$10$ auto $1.5\times 10^{-6}$ 0.00230 0.00916 0.00282 $34$ $1.0\times 10^{6}$ $2.00\times 10^{-6}$ $5.11\times 10^{-2}$ 59 58 2
-$11$ auto $1.5\times 10^{-7}$ 0.00117 0.00847 0.00269 $38$ $1.1\times 10^{7}$ $2.16\times 10^{-6}$ $5.04\times 10^{-2}$ 73 76 2
+$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+$5$ 0.00178 0.00606 $16$ $5.0\times 10^{6}$ $4.91\times 10^{-7}$ $2.84\times 10^{-1}$ 16 11 2
+$6$ 0.00154 0.00503 $20$ $6.0\times 10^{6}$ $6.37\times 10^{-7}$ $3.38\times 10^{-1}$ 16 11 2
+$7$ 0.00134 0.00428 $24$ $7.0\times 10^{6}$ $9.12\times 10^{-7}$ $3.71\times 10^{-1}$ 17 11 2
+$8$ 0.00136 0.00372 $28$ $8.0\times 10^{6}$ $1.75\times 10^{-6}$ $4.09\times 10^{-1}$ 18 11 2
+$9$ 0.000867 0.00329 $32$ $9.0\times 10^{6}$ $2.47\times 10^{-6}$ $4.51\times 10^{-1}$ 20 11 2
+$10$ 0.00229 0.00294 $10$ $1.0\times 10^{7}$ $1.18\times 10^{-6}$ $5.11\times 10^{-1}$ 20 11 1
+$11$ 0.000554 0.00265 $11$ $1.1\times 10^{7}$ $1.16\times 10^{-6}$ $5.04\times 10^{-1}$ 20 11 1
 (b) $\vec{S}_{0}$
-$d$ $\eta$ $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ $e_{\mathrm{TT}}$ $e_{\text{MC},10^{5}}$ $e_{\text{MC},10^{6}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC},10^{5}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC},10^{5}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
-$5$ 0.9 $1.0\times 10^{-5}$ 0.0195 0.0420 0.0133 $14$ $5.0\times 10^{5}$ $7.67\times 10^{-7}$ $3.99\times 10^{-2}$ 14 9 2
-$6$ 0.75 $1.0\times 10^{-5}$ 0.0165 0.0385 0.0121 $18$ $6.0\times 10^{5}$ $9.34\times 10^{-7}$ $5.15\times 10^{-2}$ 16 9 2
-$7$ 0.69 $3.0\times 10^{-5}$ 0.0103 0.0357 0.0112 $22$ $7.0\times 10^{5}$ $9.21\times 10^{-7}$ $5.37\times 10^{-2}$ $21$ $10$ $2$
-$8$ 0.57 $3.5\times 10^{-5}$ 0.0125 0.0303 0.00952 $26$ $8.0\times 10^{5}$ $1.63\times 10^{-6}$ $6.18\times 10^{-2}$ $17$ $7$ $2$
-$9$ 0.53 $1.0\times 10^{-5}$ 0.0155 0.0204 0.00650 $30$ $9.0\times 10^{5}$ $1.19\times 10^{-6}$ $7.95\times 10^{-2}$ $19$ $7$ $2$
-$10$ 0.5 $3.0\times 10^{-5}$ 0.0172 0.0204 0.00654 $34$ $1.0\times 10^{6}$ $1.47\times 10^{-6}$ $8.77\times 10^{-2}$ $16$ $6$ $2$
-$11$ auto $1.0\times 10^{-7}$ 0.0139 0.0181 0.00579 $38$ $1.1\times 10^{6}$ $1.93\times 10^{-6}$ $8.43\times 10^{-2}$ $48$ $38$ $2$
+$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+$5$ 0.00151 0.00773 $16$ $5.0\times 10^{6}$ $7.67\times 10^{-7}$ $3.99\times 10^{-1}$ 18 11 2
+$6$ 0.00122 0.00640 $20$ $6.0\times 10^{6}$ $6.34\times 10^{-7}$ $5.15\times 10^{-1}$ 19 11 2
+$7$ 0.00112 0.00547 $24$ $7.0\times 10^{6}$ $9.21\times 10^{-7}$ $5.37\times 10^{-1}$ $21$ $10$ $2$
+$8$ 0.000973 0.00477 $28$ $8.0\times 10^{6}$ $1.63\times 10^{-6}$ $6.18\times 10^{-1}$ $23$ $11$ $2$
+$9$ 0.000686 0.00424 $32$ $9.0\times 10^{6}$ $1.19\times 10^{-6}$ $7.95\times 10^{-1}$ $24$ $11$ $2$
+$10$ 0.000662 0.00377 $36$ $1.0\times 10^{7}$ $1.47\times 10^{-6}$ $8.77\times 10^{-1}$ $24$ $11$ $2$
+$11$ 0.00114 0.00339 $40$ $1.1\times 10^{7}$ $1.93\times 10^{-6}$ $8.43\times 10^{-1}$ $25$ $13$ $2$
 Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
-Here, we set the ranges $\sigma_{m}\in[0.15,0.25]$ and $S_{m,0}\in[90,120]$, and place 100 equally spaced grid points within these ranges.
-The $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ in the table refers to the values adjusted so that $c_{\mathrm{TT}}$ is reduced with $e_{\mathrm{TT}}$ kept smaller than the MC error $e_{\text{MC},10^{5}}$.
-The term auto in (a) and (b) indicate that $\eta_{m}$ was automatically determined for each asset based on Eqs.
-(19) and (20), respectively.
-The computational complexity and time increases when $\eta$ is set by Eq.
-(19) or (20).
+Here, we set the ranges $\sigma_{m}\in[0.15,0.25)$ and $S_{m,0}\in[90,120)$, and place 100 equally spaced grid points within these ranges.
 $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
 Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$ versus the number of assets $d$.
-We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{5}$ and $10^{6}$.
+We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{6}$.
 Figure 3: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$.
 The odd bond $l=2i-1$ connects the core tensors for $j_{m}$ and $k_{m}$, the indexes for $z_{m}$ and $p_{m}$, respectively, and the even bond $l=2i$ connects those for $k_{m}$ and $j_{m+1}$.
-In the cases corresponding to the left panels and the bottom right panel, $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{m}$ and $z_{m+1}$).
+Note that $\hat{v}$ does not depend on parameters, and thus the bond $l$ takes only even values $2i$ (connecting the core tensors for $z_{m}$ and $z_{m+1}$).
 It is noteworthy that the graph of bond dimensions exhibits a characteristic jagged shape.
@@
 3(b)).
-For $d=11$, when $\eta$ is determined automatically, this bond dimension increases to around 40.
-The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 2 Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
-Compared with the MC-based method with $10^{5}$ paths, the TT-based mathod again gives the smaller error in every value of $d$.
-With $10^{6}$ paths, the TT error is worse than the MC error for many values of $d$, but still comparable.
+The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 2.
+Therefore, we saw superiority over the Monte Carlo method in terms of computation complexity and time.
+Compared with the MC-based method with $10^{6}$ paths, the TT-based mathod again gives the smaller error in every value of $d$.
 VI.0.3 Randomness in learning the TTs
 Here, we mention the randomness in learning the TTs and the error induced by it.
-Since TCI starts with the randomly selected grid points, the accuracy of the TT-based method depends on these initial points.
-To assess such a fluctuation of the accuracy, for $d=10$, we evaluated the mean and standard deviation of the relative accuracy of the TT-based method in 20 runs, with initial points randomly selected in each.
-As a result, the mean was 0.0324 and the standard deviation was 0.0039, which indicates that the accuracy fluctuation of TT-based option pricing is very small.
 We should note that TCI is a heuristic method, and depending on the choice of the initial points, the learning might not work well.
@@
 (3) might not go below the threshold.
-Here, with the tolerance for TCI/SVD fixed, if the maximum relative error exceeds 4%, we redo TCI and SVD.
-The aforementioned mean and standard deviation are computed from successful cases.
+To assess such a fluctuation of the accuracy, for $d=10$ in the case of varying $\vec{\sigma}$, we evaluated the mean and standard deviation of the accuracy of the TT-based method in 20 runs, with initial points randomly selected in each.
+As a result, the mean was 0.00230 and the standard deviation was $4.74\times 10^{-9}$, which indicates that the accuracy fluctuation of TT-based option pricing is very small.
 VI.0.4 Total computational time for obtaining the TTs
@@
 We focus on TCI because it dominates over SVD.
-It took about 3.4 hours for $d=11$ in the case that $\vec{\sigma}$ dependence is involved.
+It took about $12.5$ minutes for $d=11$ in the case that $\vec{S}_{0}$ dependence is involved.
 This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec.
 VII.
-We also note that, when we do not incorporate the parameter dependence into the tensor trains for $\phi$ and $\hat{v}$ as in Ref.
+We also note that, when we do not incorporate the parameter dependence into the tensor train for $\phi$ as in Ref.
 [6], TCI takes a much longer time than the MC-based method, 7.3 seconds for $d=11$ in the case of $\vec{\sigma}$ dependence.
-In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$ and all other parameters are the same as those in the default settings listed in Table 1.
+In this setup, we set $\eta=0.48$, $\epsilon_{\mathrm{TCI}}=10^{-6}$, $N=50$ and all other parameters are the same as those in the default settings listed in Table 1.
 This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
 VII Discussion
-We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from functions of Eqs.
-(10) and (12), enabling fast option pricing in response to varying parameters.
+We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from the function of Eq.
+(10), enabling fast option pricing in response to varying parameters.
 In this study, we considered scenarios with varying volatility and present stock prices as benchmarks for our proposed method.
 Up to $d=11$, we demonstrated superiority in both computational complexity and time.
-Note that in the MC-based method, the implementation may have room for improvement to reduce computational time.
-We numerically verified that all TT-based results fell within the statistical error range of the MC-based method with $10^{5}$ paths.
-Compared to the MC-based method with $10^{6}$ paths, the TT-based method gave the worse errors in some cases, but was still comparable.
-To enhance the accuracy of the TT-based method, we may reduce the tolerance of the TCI further or change the initial randomly selected grid points of TCI.
+Note that in the MC-based method, the implementation may have room for improvement to reduce computational complexity and time.
+We have also seen that in all the tested cases, the error in the TT-based result is smaller than that of the MC-based result with $10^{6}$ paths.
 Now, let us consider how the proposed method provides benefits in practical business in financial firms.
 An expected way to utilize this method is as follows.
-At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the TT to quickly price the option for the fluctuating parameters.
+At night, when the financial market is closed, we learn the TTs and perform contractions, and then, during the day, when the market is open, we use the TT of parameter-dependent option prices to quickly price the option for the fluctuating parameters.
 If we pursue the computational speed in the daytime and allow the overnight precomputation to some extent, the above operation can be beneficial.
@@
 Finally, we discuss future research directions.
-If we could keep $\eta$ constant when $d$ is increased, we have the advantage of reducing the computational complexity involved in option pricing.
-It may be possible to keep $\eta$ constant by adjusting the value of parameter $N$, e.g.
-by increasing $N$ from the current value of 50 to, say, 100.
-It could be a promising approach to directly learning the TTOs incorporating parameter dependence using TCI, potentially improving the accuracy of option pricing.
-At the same time, it is important to study methods to improve the numerical stability and accuracy of TCI itself.
 For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step.
 Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing.
-With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [38] and stochastic volatility model [39, 40] are also used in practice.
+With respect to the pricing model, while this paper has considered the BS model, more advanced models such as the local volatility model [39], stochastic volatility model [40, 41] and Lévy models [42, 43] are also used in practice.
 Applying our method to such models, where the dependencies on the parameters in the models are incorporated into TTs, is an interesting direction.
@@
 Considering whether our method can be extended for such options will be an important and interesting challenge.
-VIII Acknowledgements
+Acknowledgements
 We are grateful to Marc K.
@@
 This work was supported by the Center of Innovation for Sustainable Quantum AI (JST Grant Number JPMJPF2221).
-IX Author contributions statement
-R.S.
-conceived the project, and R.S.
-and H.T.
-wrote the code with help from K.M.
-and performed numerical simulations.
-All authors actively discussed the results and contributed to the preparation of the manuscript.
-X Appendix
-X.1 How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
-Figure 4: The maximum error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
-Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
-Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-9}$.
-The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$.
+Appendix A How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
+Figure 4: The mean absolute error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs when incorporating the $\vec{\sigma}$ dependence at $d=10$.
+Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ to the TT when incorporating $\vec{\sigma}$ dependence at $d=10$ obtained through TCI.
+Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-6}$.
+The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$.
 From FIG.
-4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$.
+4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-2}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $1.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$.
 By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
-From the fact that we can keep the maximum relative error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information.
+From the fact that we can keep the error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information.
 By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed.
-In addition, it is surprising that in the analysis with $\epsilon_{\text{SVD}}=10^{-6}$, the maximum relative error decreased compared to before compression by SVD.
+In addition, it is surprising that in the analysis with $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ = $10^{-6}$, the $e_{\mathrm{TT}}$ decreased compared to before compression by SVD.
 We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance.

### v8

@@
 As a benchmark test, we run the proposed method to price a multi-asset option for the various values of volatilities and present asset prices.
-We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{6}$ paths in terms of computational complexity while keeping better accuracy.
+We show that, in the tested cases involving up to 11 assets, the proposed method outperforms Monte Carlo-based option pricing with $10^{5}$ and $10^{6}$ paths in terms of computational complexity while keeping comparable accuracy.
 I Introduction
@@
 Concerning the time when the payoff occurs, the simplest and most common type of options is the European type, which this paper hereafter focuses on: at the predetermined future time (maturity) $T$, the payoff $v(\vec{S}(T))$ depending on the underlying asset prices $\vec{S}(T)$ at time $T$ occurs.
-For example, in a European call/put option, one party has the right to buy/sell an asset at the predetermined price (strike) $K$ and maturity $T$, and the corresponding payoff function is $v(S(T))=\max\{c(S(T)-K),0\}$, where $c=1$ and $-1$ for a call and put option, respectively.
+For example, in a single-asset European call/put option, one party has the right to buy/sell an asset at the predetermined price (strike) $K$ and maturity $T$, and the corresponding payoff function is $v(S(T))=\max\{c(S(T)-K),0\}$, where $c=1$ and $-1$ for a call and put option, respectively.
 In addition to this simple one, various types of options are traded and constitute a large part of the financial industry.
@@
 In particular, pricing multi-asset options, whose payoff depends on the prices of multiple underlying assets, is often demanding.
-Many pricing methodologies suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
-The quasi-Monte Carlo method has a better convergence rate in low-dimensional cases, but it can deteriorate in high-dimensional cases.
+Pricing methodologies such as PDE-based solvers (e.g., the finite difference method [3]) and tree models [4] suffer from the so-called curse of dimensionality, which means the exponential increase of computational complexity with respect to the asset number, and the Monte Carlo method, which may evade the exponential complexity, has a slow convergence rate.
+The quasi-Monte Carlo method is more effective than the Monte Carlo method up to a certain level of the dimension but has the worst-case complexity exponential with respect to the dimension, despite the methods for mitigating this proposed [5, 6, 7].
+At any rate, in the rapidly changing financial market, the pursuit of the speed-up of option pricing can never go too far.
 Motivated by these points, recently, applications of quantum computing to option pricing are considered actively [^2] See Ref.
@@
 In light of this, applications of quantum-inspired classical algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present  [10, 11, 12, 13].
-Among them, this paper focuses on the application of tensor train (TT) [9]  [^3] Tensor train is a mathematically equivalent to matrix product state (MPS) [10].
-learning to the Fourier transform (FT)-based option pricing method [11, 12], following the original proposal in Ref.
+Among them, this paper focuses on the application of tensor train (TT) [14]learning to the FT-based option pricing method, following the original proposal in Ref.
 [11].
@@
 After this conversion, the numerical integration is done more efficiently in many cases.
-Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially.
+Unfortunately, the FT-based method with naive grid-based integration suffers from the curse of dimensionality, with its computational complexity increasing exponentially for multi-asset options.
 On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [15, 16] as reviews.
-and recently, it has also been utilized in various fields such as machine learning [15, 16, 17], quantum field theory [18, 19, 20], and partial differential equations [7, 21, 22, 23, 24].
+and recently, it has also been utilized in various fields such as machine learning [17, 18, 19], quantum field theory [20, 21, 22], and partial differential equations [12, 23, 24, 25].
 Ref.
@@
 Here, we would like to point out an issue in this TT-based method.
-That is, we need to rerun TCI to obtain TTs each time the input parameters of option price are changed.
-According to our numerical experiments, TCI takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
+That is, we need to rerun TCI to obtain TTs and compute option price each time the input parameters such as volatility and initial asset price, are changed.
+According to our numerical experiments, the TT learning method for FT-based option pricing [11] takes a longer time than the Monte Carlo method, which means that the TT-based method does not have a computational time advantage.
 To address this issue, we focus on another usage of tensor trains that can embed parameter dependence [30, 31, 32, 33] to make FT-based option pricing more efficient.
@@
 1).
-To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices.
+Note that this is an advantage of the proposed method not only over the original TT-bsed method in Ref.
+[11] but also over other aforementioned methods such as PDE-based solvers, tree models, and the (quasi-)Monte Carlo method.
+When parameters are changed and an option is repriced, these methods must be rerun from scratch, in contrast to the proposed method, in which we just reuse the TT that is evaluated efficiently.
+For evaluation, we consider two benchmark scenarios in which we vary volatilities and present stock prices under the Black–Scholes model, focusing specifically on a min-option.
 In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{6}$ paths by a factor of $O(10^{5})$ (see Fig.
-2).
+LABEL:fig:table_cc).
 We also confirm numerically that in the tested cases, the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{6}$ paths.
-In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the better accuracy.
+In summary, these results indicate that, at least in the tested cases, our proposed method offers significant advantages in computational complexity while keeping a better accuracy.
 In the context of TT-based approximations of high-dimensional functions, incorporating parameter dependence into TTs has been considered in some fields [30, 31, 32, 33].
 However, to the best of our knowledge, our study is the first to take such an approach in FT-based option pricing in order to make it more efficient for varying parameters, which provides practical benefits in the rapidly changing financial market.
-The rest of this paper is organized as follows.
-Sec.
-II is devoted to introducing the tensor train and tensor learning algorithm.
-In Sec.
-III, we review the FT-based option pricing and its reformulation aided by TCI.
-We propose a new scheme for fast option pricing by learning tensor trains with parameter dependence in Sec.
-IV.
-We show the results of the numerical demonstration of our method applied to a kind of multi-asset option for the various values of the volatilities and the present asset prices in Sec.
-V.
-The summary and discussions are given in Sec.
-VI.
 II Tensor train
@@
 $F_{x_{1},\ldots,x_{d}}$ $\approx\sum_{l_{1}}^{\chi_{1}}\cdots\sum_{l_{d-1}}^{\chi_{d}}F_{l_{1},x_{1}}^{(1)}F_{l_{1}l_{2},x_{2}}^{(2)}\cdots F_{l_{d-1},x_{d}}^{(d)}$
-$\equiv F_{x_{1}}^{(1)}\cdot F_{x_{2}}^{(2)}\cdot(\cdots)\cdot F_{x_{d}}^{(d)}$ (1)
+$\equiv\prod_{i=1}^{d}F_{x_{i}}^{(i)}$ (1)
 where $F_{x_{i}}^{(i)}$ denotes each 3-way tensor, $l_{i}$ represents the virtual bond index, and $\chi_{i}$ is the dimension of the virtual bond.
 One of the main advantages of TT is that it significantly reduces computational complexity and memory requirements by reducing bond dimensions $\chi_{i}$.
+Tensor train is a mathematically equivalent to matrix product state (MPS) [34].
 This is an equivalent expression to the wave function $F_{x_{1},\ldots,x_{d}}$ of a quantum system with $d$ $N$-level qudits as follows:
@@
 […]
+In the same manner, the $2d$-way tensor $F_{x_{1},\dots,x_{d}}^{y_{1},\dots,y_{d}}$ where each local index $x_{l}$ has local dimension $N_{l}$ and each local index $y_{l}$ has local dimension $M_{l}$, $l=1,\dots,d$, can be expressed with the product of the each tensor core, i.e., fourth-order tensor $[F_{a_{i-1}x_{i}a_{i}}^{y_{i}}]^{(i)}$,
+$F_{x_{1},\ldots,x_{d}}^{y_{1},\ldots,y_{d}}$ $\approx\sum_{a_{1}=1}^{\chi_{1}}\cdots\sum_{a_{d-1}=1}^{\chi_{d}}[F_{a_{0}x_{1}a_{1}}^{y_{1}}]^{(1)}\cdots[F_{a_{d-1}x_{d}a_{d}}^{y_{d}}]^{(d)}$
+$=\prod_{i=1}^{d}[F_{x_{i}}^{y_{i}}]^{(i)}$ (3)
+is called the tensor train operator (TTO) or matrix product operator (MPO) [34].
 II.1 Compression tenchniques
@@
 $\vec{z}_{\mathrm{gr},\vec{j}}:=(\eta j_{1},\ldots,\eta j_{d}),$ (15)
-where $\eta$ is the integration step size in each direction, and $\Delta_{\rm vol}:=\eta^{d}$ is the volume element.
+where $\eta$ is the constant integration step size in each direction, and $\Delta_{\rm vol}:=\eta^{d}$ is the volume element.
 $\eta$ is a hyperparameter that must be appropriately determined.
-In our demonstration, it is set to 0.4 or 0.3, which yields the accurate result (see Sec.
-V for details).
+In our demonstration, it is set to 0.4 or 0.3, which yields the accurate result (see Numerical demonstration for details).
 III.2 Fourier transform-based option pricing with tensor trains
@@
 (14) by
-$V(\vec{\sigma},\vec{S}_{0})\simeq\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\tilde{\phi}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}}\Delta_{\mathrm{vol}}.$ (15)
+$V(\vec{p})\simeq\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\tilde{\phi}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}}\Delta_{\mathrm{vol}},$ (16)
+where each index $j_{i}$ ($i=1,\cdots,d$) is from $-\frac{N}{2}$ to $\frac{N}{2}$ and the corresponding integration range for each variable is $-\frac{\eta N}{2}$ to $\frac{\eta N}{2}$.
+We must suitably choose both the number of grid points $N$ and the integral step size $\eta$, which serve as hyperparameters in this approach.
 Thanks to TCI, we can obtain the approximate TTs avoiding the evaluations of $\phi$ and $\hat{v}$ at all the grid points.
@@
 Then, as illustrated in FIG.
-1 (a), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},\ldots,j_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ and $(\vec{z}_{\mathrm{gr},\vec{j}})$, respectively.
+1 (a) (1), we run TCI to get the TTs $\tilde{\phi}_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\tilde{v}_{j_{1},\ldots,j_{d}}$ that respectively approximate the tensors $\phi_{j_{1},k_{1},\ldots,j_{d},k_{d}}$ and $\hat{v}_{j_{1},\ldots,j_{d}}$, whose entries are the values of $\phi$ and $\hat{v}$ at grid points $(\vec{z}_{\mathrm{gr},\vec{j}},\vec{p}_{\mathrm{gr},\vec{k}})$ and $(\vec{z}_{\mathrm{gr},\vec{j}})$, respectively.
 We reduce the bond dimensions of these TTs using SVD.
-We then contract adjacent core tensors pairwise and obtain a tensor train operators (TTO), also known as matrix product operators (MPO), $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$.
-We contract this TTO $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$.
-and TT $\tilde{v}_{j_{1},\ldots,j_{d}}$ along the index vector $\vec{j}$ as follows:
+As shown in FIG.
+1 (a) (2), we then contract adjacent core tensors pairwise and obtain a tensor train operators (TTO) $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$.
+Then, As shown in FIG.
+1 (a) (3), we contract this TTO $\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}$ and TT $\tilde{v}_{j_{1},\ldots,j_{d}}$ along the index vector $\vec{j}$ as follows:
 $\tilde{V}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}_{j_{1},\ldots,j_{d}},$ (18)
-Then we optimize the bond dimensions of $\tilde{V}_{k_{1},\ldots,k_{d}}$ using SVD.
-Thus, we get a new function $\tilde{V}_{k_{1},\ldots,k_{d}}$ whose input are parameter and output is option price.
+Then, we optimize the bond dimensions of $\tilde{V}_{k_{1},\ldots,k_{d}}$ using SVD.
+Thus, as shown in FIG.
+1 (a) (4), we get a new function $\tilde{V}_{k_{1},\ldots,k_{d}}$ whose input are parameter and output is option price.
 Having this TT, we get the option price for the specified value of $\vec{p}$ as illustrated in FIG.
@@
 Two core tensors in the TT that correspond to $z_{j}$ and $p_{j}$ of the same asset are arranged next to each other, i.e., the order is ($z_{1}p_{1}z_{2}p_{2}\cdots z_{d}p_{d}$).
-In the demonstration in Sec.
-V, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing.
+In the numerical demonstration, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing.
 On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains.
@@
 Although we do not discuss it in detail, this is an important topic for future research.
-In the two test cases for our proposed method in Sec.
-V, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices.
-Note that it is possible to include dependencies on other parameters in the TTs.
-For example, taking into account the dependence on the parameters concerning the option contract, such as the maturity $T$ and the strike $K$, enables us to price different option contracts with a single set of TTs.
-Although this is a promising approach, there might be some issues.
-For example, it is non-trivial whether TTs incorporating many parameter dependencies have a low-rank structure.
-Thus, we will leave such a study for future work.
+In the two test cases for our proposed method in Numerical demonstration, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices.
 Figure 1: Fast option pricing based on TNs proposed in this study.
@@
 Here, we denote the maximum bond dimensions as $\chi_{V}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations.
-Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec.
-V.
+Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Discussion.
 Here, we ignore the computational complexity of all the processes in FIG.
@@
 This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing.
-As discussed in Sec.
-VII, we can reasonably find such a situation in practice.
+As discussed in Discussion, we can reasonably find such a situation in practice.
 V Numerical demonstration
@@
 In the following, we describe the parameter values used in this study, the software used, and how the errors were evaluated.
-Ranges of ${\sigma_{m}}$ and ${S_{m,0}}$
+V.1.1 Ranges of volatility and initial stock price
 With respect to $\vec{p}$, on which the TTs learn the dependence of the functions $\phi$ and $\hat{v}$, we take the two test cases: $\vec{p}$ is the volatilities $\vec{\sigma}$ or the present asset prices $\vec{S}_{0}$.
@@
 For both $\sigma_{m}$ and $S_{m,0}$, we set 100 equally spaced grid points in the range, and so the total number of the grid points in the space of $\vec{p}$ is $100^{d}$.
-Other parameters
+V.1.2 Other parameters
 The other parameters for option pricing are fixed to the values summarized in the table below.
-$T$ $r$ $K$ $S_{0}$ $\alpha$ $\rho_{mn},m\neq n$ $(N,\eta)$
-$1$ $0.01$ $100$ $100$ $\frac{5}{d}$ $\frac{1}{3}$$\left\{\begin{array}[]{ll}(200,0.3)&;\ d=11,\vec{p}=\vec{S}_{0}\\ (100,0.4)&;\ \mathrm{otherwise}\end{array}\right.$
-Table 1: The parameters except $\vec{\sigma}$ and $\vec{S}_{0}$.
-Error evaluation
+The values of $\alpha$ are adopted from previous study [11] and we confirmed that the solution remains stable when the values are varied slightly.
+$N$ and $\eta$ are set sufficiently large and small, respectively, so that the accuracy of the proposed approach becomes better than the MC.
+$T$ $r$ $K$ $S_{0}$ $\sigma$ $\alpha$ $\rho_{mn}\ (m\neq n)$ $(N,\eta)$
+$1$ $0.01$ $100$ $100$ $0.2$ $\frac{5}{d}$ $\frac{1}{3}$$\begin{cases}(200,0.3)&\text{if }d=11,\ \vec{p}=\vec{S}_{0}\\ (100,0.4)&\text{otherwise}\end{cases}$
+Table 1: The input parameters used in the experiment, except for $\vec{\sigma}$ and $\vec{S}_{0}$.
+When either $\vec{\sigma}$ or $\vec{S}_{0}$ is selected as the parameter dependency $\vec{p}$, the remaining one is fixed at the value shown in the table.
+For the case where $\vec{S}_{0}$ is varied and $d=11$, we use $(N,\eta)=(200,0.3)$.
+In all other cases, we use $(N,\eta)=(100,0.4)$.
+V.1.3 Error evaluation
 We do not have the exact price of the multivariate min-call option since there is no known analytic formula for it.
@@
 We compare the mean absolute error of our method for the 100 parameter sets with the one obtained from the Monte Carlo simulations with the same parameter setting.
-Software and hardware used in this study
+V.1.4 Software and hardware used in this study
 TensorCrossInterpolation.jl [38] was used for learning tensor trains from functions appearing Fourier-based option pricing.
-ITensors.jl [37] was utilized for computing the contraction and inner product of the tensor trains.
 The Monte Carlo simulations were carried out using tf-quant-finance [39].
@@
 The computations were performed on a 2023 MacBook Pro featuring a 12-core Apple M2 Max processor and 32GB of 400GB/s unified memory.
-VI Results
+V.2 Results
 We show the results for the computational complexity, time, and accuracy of TT-based and MC-based option pricing when two parameters $\vec{\sigma}$ and $\vec{S_{0}}$ are varied.
+(a) $\vec{\sigma}$
+$d$ $e_{\mathrm{TT}}$ $e_{\mathrm{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}\,[$s]]]] $t_{\mathrm{MC}}\,[$s]]]] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+5 0.00178 0.00606 16 $5.0\times 10^{6}$ $4.91\times 10^{-7}$ $2.84\times 10^{-1}$ 16 11 2
+6 0.00154 0.00503 20 $6.0\times 10^{6}$ $6.37\times 10^{-7}$ $3.38\times 10^{-1}$ 16 11 2
+7 0.00134 0.00428 24 $7.0\times 10^{6}$ $9.12\times 10^{-7}$ $3.71\times 10^{-1}$ 17 11 2
+8 0.00136 0.00372 28 $8.0\times 10^{6}$ $1.75\times 10^{-6}$ $4.09\times 10^{-1}$ 18 11 2
+9 0.000867 0.00329 32 $9.0\times 10^{6}$ $2.47\times 10^{-6}$ $4.51\times 10^{-1}$ 20 11 2
+10 0.00229 0.00294 10 $1.0\times 10^{7}$ $1.18\times 10^{-6}$ $5.11\times 10^{-1}$ 20 11 1
+11 0.000554 0.00265 11 $1.1\times 10^{7}$ $1.16\times 10^{-6}$ $5.04\times 10^{-1}$ 20 11 1
+(b) $\vec{S}_{0}$
+$d$ $e_{\mathrm{TT}}$ $e_{\mathrm{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}\,[$s]]]] $t_{\mathrm{MC}}\,[$s]]]] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
+5 0.00151 0.00773 16 $5.0\times 10^{6}$ $7.67\times 10^{-7}$ $3.99\times 10^{-1}$ 18 11 2
+6 0.00122 0.00640 20 $6.0\times 10^{6}$ $6.34\times 10^{-7}$ $5.15\times 10^{-1}$ 19 11 2
+7 0.00112 0.00547 24 $7.0\times 10^{6}$ $9.21\times 10^{-7}$ $5.37\times 10^{-1}$ 21 10 2
+8 0.000973 0.00477 28 $8.0\times 10^{6}$ $1.63\times 10^{-6}$ $6.18\times 10^{-1}$ 23 11 2
+9 0.000686 0.00424 32 $9.0\times 10^{6}$ $1.19\times 10^{-6}$ $7.95\times 10^{-1}$ 24 11 2
+10 0.000662 0.00377 36 $1.0\times 10^{7}$ $1.47\times 10^{-6}$ $8.77\times 10^{-1}$ 24 11 2
+11 0.00114 0.00339 40 $1.1\times 10^{7}$ $1.93\times 10^{-6}$ $8.43\times 10^{-1}$ 25 13 2
+Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
+Here, we set the ranges $\sigma_{m}\in[0.15,0.25)$ and $S_{m,0}\in[90,120)$, placing 100 equally spaced grid points within each range.
+$t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ are averages over 100 measurements.
 The results are summarized in TABLE 2.
 In particular, the computational complexity versus $d$ is plotted in FIG.
-2.
+LABEL:fig:table_cc.
 The mean absolute error in TT-based option prices among runs for 100 random parameter sets is represented by $e_{\mathrm{TT}}$.
@@
 In the SVD on $\tilde{V}_{k_{1},\ldots,k_{d}}$ obtained by contracting the TTs for $\phi$ and $\hat{v}$, we set the tolerance $\epsilon^{\tilde{V}}_{\mathrm{SVD}}=1.0\times 10^{-6}$, and the maximum bond dimension of the resulting TT is denoted by $\chi_{V}$.
-VI.0.1 The case of varying $\vec{\sigma}$
+V.2.1 The case of varying volatilities
 TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence.
@@
 The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in FIG.
-3(a).
-By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10 when $\eta$ is constant.
-The details of this compression by SVD are described in Appendix.
-A.
+2(a).
+By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at around 10.
+The details of this compression by SVD are described in A.
 The maximum bond dimension $\chi_{\tilde{V}}$ of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is maintained at 2, or 1 especially when $d=10,11$.
@@
 Besides, the error in the TT-based result is smaller than that in the MC-based result in all the tested cases.
-(a) $\vec{\sigma}$
-$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
-$5$ 0.00178 0.00606 $16$ $5.0\times 10^{6}$ $4.91\times 10^{-7}$ $2.84\times 10^{-1}$ 16 11 2
-$6$ 0.00154 0.00503 $20$ $6.0\times 10^{6}$ $6.37\times 10^{-7}$ $3.38\times 10^{-1}$ 16 11 2
-$7$ 0.00134 0.00428 $24$ $7.0\times 10^{6}$ $9.12\times 10^{-7}$ $3.71\times 10^{-1}$ 17 11 2
-$8$ 0.00136 0.00372 $28$ $8.0\times 10^{6}$ $1.75\times 10^{-6}$ $4.09\times 10^{-1}$ 18 11 2
-$9$ 0.000867 0.00329 $32$ $9.0\times 10^{6}$ $2.47\times 10^{-6}$ $4.51\times 10^{-1}$ 20 11 2
-$10$ 0.00229 0.00294 $10$ $1.0\times 10^{7}$ $1.18\times 10^{-6}$ $5.11\times 10^{-1}$ 20 11 1
-$11$ 0.000554 0.00265 $11$ $1.1\times 10^{7}$ $1.16\times 10^{-6}$ $5.04\times 10^{-1}$ 20 11 1
-(b) $\vec{S}_{0}$
-$d$ $e_{\mathrm{TT}}$ $e_{\text{MC}}$ $c_{\mathrm{TT}}$ $c_{\mathrm{MC}}$ $t_{\mathrm{TT}}$[s] $t_{\mathrm{MC}}$[s] $\chi_{\phi}$ $\chi_{\hat{v}}$ $\chi_{\tilde{V}}$
-$5$ 0.00151 0.00773 $16$ $5.0\times 10^{6}$ $7.67\times 10^{-7}$ $3.99\times 10^{-1}$ 18 11 2
-$6$ 0.00122 0.00640 $20$ $6.0\times 10^{6}$ $6.34\times 10^{-7}$ $5.15\times 10^{-1}$ 19 11 2
-$7$ 0.00112 0.00547 $24$ $7.0\times 10^{6}$ $9.21\times 10^{-7}$ $5.37\times 10^{-1}$ $21$ $10$ $2$
-$8$ 0.000973 0.00477 $28$ $8.0\times 10^{6}$ $1.63\times 10^{-6}$ $6.18\times 10^{-1}$ $23$ $11$ $2$
-$9$ 0.000686 0.00424 $32$ $9.0\times 10^{6}$ $1.19\times 10^{-6}$ $7.95\times 10^{-1}$ $24$ $11$ $2$
-$10$ 0.000662 0.00377 $36$ $1.0\times 10^{7}$ $1.47\times 10^{-6}$ $8.77\times 10^{-1}$ $24$ $11$ $2$
-$11$ 0.00114 0.00339 $40$ $1.1\times 10^{7}$ $1.93\times 10^{-6}$ $8.43\times 10^{-1}$ $25$ $13$ $2$
-Table 2: The results of TT-based option pricing incorporating (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$ dependence.
-Here, we set the ranges $\sigma_{m}\in[0.15,0.25)$ and $S_{m,0}\in[90,120)$, and place 100 equally spaced grid points within these ranges.
-$t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$ represent the average computational time from 100 measurements, respectively.
-Figure 2: The computational complexity of TT-based and MC-based option pricing, denoted as $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$ versus the number of assets $d$.
-We consider including parameter dependencies on $\vec{\sigma}$ and $\vec{S_{0}}$ in TTs and the number of path in MC are $10^{6}$.
 Figure 2: The bond dimensions $\chi_{l}$ of each bond $l$ for $\phi$ and $\hat{v}$, obtained through TCI and SVD, incorporating dependencies on (a) $\vec{\sigma}$ and (b) $\vec{S}_{0}$.
@@
 The bond dimensions between sites for $j_{m}$ and $k_{m}$, which are related to the same asset, are large, and the ones between sites for $k_{m}$ and $j_{m+1}$, which are related to different assets, are small.
-VI.0.2 The case of varying $\vec{S_{0}}$
+V.2.2 The case of varying initial stock prices
 TABLE 2 (b) shows the results of TT-based option pricing when we consider $\vec{S_{0}}$ dependence.
@@
 For $d$=5 and 10, after compression using SVD and contractions, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ were reduced to around 10 (refer to FIG.
-3(b)).
+2(b)).
 The maximum bond dimension, $\chi_{\tilde{V}}$, of $\tilde{V}_{k_{1},\ldots,k_{d}}$ is 2.
@@
 Compared with the MC-based method with $10^{6}$ paths, the TT-based mathod again gives the smaller error in every value of $d$.
-VI.0.3 Randomness in learning the TTs
+V.2.3 Randomness in learning the TTs
 Here, we mention the randomness in learning the TTs and the error induced by it.
@@
 As a result, the mean was 0.00230 and the standard deviation was $4.74\times 10^{-9}$, which indicates that the accuracy fluctuation of TT-based option pricing is very small.
-VI.0.4 Total computational time for obtaining the TTs
+V.2.4 Total computational time for obtaining the TTs
 We mention the total computational time for obtaining the TTs for $\phi$ and $\hat{v}$ and $\tilde{V}_{k_{1},\ldots,k_{d}}$ through TCI and SVD.
@@
 It took about $12.5$ minutes for $d=11$ in the case that $\vec{S}_{0}$ dependence is involved.
-This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Sec.
-VII.
+This is sufficiently short for a practical use-case of the TT-based method mentioned exemplified in Discussion.
 We also note that, when we do not incorporate the parameter dependence into the tensor train for $\phi$ as in Ref.
@@
 This means that running TCI every time the parameter varies does not lead to the time advantage of the TT-based method over the MC-based one.
-VII Discussion
+VI Discussion
 We propose a method that employs a single TCI to learn TTs incorporating parameter dependence from the function of Eq.
@@
 Finally, we discuss future research directions.
-For our method to be more practical, compressing all the input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TN format is desired, although this study tried to incorporate either of them as the first step.
+For our method to be more practical, compressing many kinds of input parameters including both $\vec{\sigma}$ and $\vec{S}_{0}$ into a single TT is desired.
+For example, taking into account the dependence on the parameters concerning the option contract, such as the maturity $T$ and the strike $K$, enables us to price different option contracts with a single set of TTs.
+Although this is a promising approach, there might be some issues.
+For example, it is non-trivial whether TTs incorporating many parameter dependencies have a low-rank structure.
+Thus, we will leave such a study for future work.
 Besides, to enhance the practical benefit, we should extend our methodology so that it is applicable to more advanced settings in option pricing.
@@
 Considering whether our method can be extended for such options will be an important and interesting challenge.
-Acknowledgements
+VII Acknowledgements
 We are grateful to Marc K.
@@
 Appendix A How to set the tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$
-Figure 4: The mean absolute error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs when incorporating the $\vec{\sigma}$ dependence at $d=10$.
-Figure 4 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ to the TT when incorporating $\vec{\sigma}$ dependence at $d=10$ obtained through TCI.
-Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-6}$.
-The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$.
+Figure 3: The maximum error $e_{\mathrm{TT}}$ when applying SVD with various $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ values to the TTs under the parameter settings of $d=10,\vec{\sigma},\eta=\text{auto}$.
+Figure 3 shows the maximum relative error $e_{\mathrm{TT}}$ when applying SVD with varying $\epsilon_{\mathrm{SVD}}$ to the TT with parameters for $d=10,~{}\vec{\sigma}$, obtained through TCI.
+Here, we fixed the tolerance of SVD $\epsilon^{\tilde{V}}_{\mathrm{SVD}}$ to $10^{-9}$.
+The tolerance $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ for SVD was chosen to maintain $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$.
 From FIG.
-4, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-2}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $1.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC}}$.
+3, it can be seen that $e_{\mathrm{TT}}$ increases sharply between $\epsilon_{\mathrm{SVD}}=10^{-6}$ and $10^{-5}$, suggesting that setting $\epsilon_{\text{SVD}}$ around $5.0\times 10^{-6}$ is appropriate for keeping $e_{\mathrm{TT}}$ smaller than $e_{\mathrm{MC},10^{5}}$.
 By compressing the bond dimension with SVD, the computational time for the contraction of these optimized TTs also can be kept low.
-From the fact that we can keep the error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information.
+From the fact that we can keep the maximum relative error $e_{\mathrm{TT}}$ small through SVD, it is suggested that the TTs obtained by TCI contain redundant information.
 By using SVD to get an optimal approximation in terms of the Frobenius norm, the reduced redundant information could be effectively removed.
-In addition, it is surprising that in the analysis with $\epsilon^{\phi,\hat{v}}_{\mathrm{SVD}}$ = $10^{-6}$, the $e_{\mathrm{TT}}$ decreased compared to before compression by SVD.
+In addition, it is surprising that in the analysis with $\epsilon_{\text{SVD}}=10^{-6}$, the maximum relative error decreased compared to before compression by SVD.
 We consider that the error contained in TTs obtained by TCI is eliminated through SVD by chance.